### v1

# Nocturnal Eye–Inspired Liquid–to-Gas Phase Change Soft Actuator with Laser-Induced-Graphene : Enhanced Environmental Light Harvesting and Photothermal Conversion

Abstract
The mobility of robotic systems is fundamentally constrained by their power sources and wiring requirements. While electrical actuation systems have achieved autonomy through battery power and wireless control, pneumatic actuators remain tethered to air supply sources. Liquid-to-gas phase change actuators utilizing low-boiling-point liquids offer a potential liquid, though they typically require substantial thermal input through heating elements that maintain electrical dependencies. External heat sources, particularly light energy, present an alternative for terrestrial applications. However, silicone-based materials, despite their optical transparency, have high volumetric heat capacity and low thermal conductivity, limiting efficient photothermal energy transfer. Previous attempts to address this through graphene or metallic powder incorporation have compromised material properties, including reduced transparency and altered elastic moduli. Inspired by the tapetum lucidum structure found in nocturnal animals’ eyes for efficient light utilization in low-light conditions, this study proposes a novel anisotropic bilayer soft actuator incorporating Laser-Induced Graphene (LIG) on the inner surface of the light-irradiated silicone layer. This creates an anisotropic structure with enhanced photothermal conversion capabilities while maintaining the advantageous properties of silicone. Comparative analysis demonstrates that the proposed actuator exhibits significantly higher photo-induced bending efficiency compared to conventional silicone-based actuators. The response time (T63) improved by 54%, decreasing from 142 seconds for pure silicone to 65 seconds, with recovery response time showing a 48% improvement. This design maintains silicone’s transparency and flexibility while utilizing LIG, which can be fabricated under ambient conditions, facilitating both manufacturing and diverse applications.
1 Introduction
Currently, most robotic actuators rely on electrical energy. For example, although pneumatic actuators use air as the driving source, they rely on electrical energy for control and to supply pressed air  (?) (?). Therefore, they cannot be considered completely independent of electrical energy. Throughout history, humans have utilized various forms of energy to enrich their lives. Aiming to use energy sources other than electrical energy to drive soft robots can significantly contribute to diversifying actuation methods in robotics.
Several methods have been proposed on using heating methods, such as laser irradiation  (?) (?), micorwave  (?) , heatsink  (?) and electromagnetic  (?) (?). Kim et al. developed an electrothermal pneumatic actuator using nanofiber mats, achieving a lightweight soft actuator that can operate without external devices by utilizing liquid evaporation  (?). However, this approach faces issues such as high power dependency and slow response times.
Various proposals have been made, starting with actuators like pouch motors ( (?))( (?))( (?))( (?))that utilize liquid-to-gas phase change, including pneumatic actuators driven by ambient temperature ( (?)). However, these actuators operate relatively slowly, with many functioning at frequencies below 0.1 Hz. This is due to significant losses in converting environmental energy into driving energy and issues related to heat transfer and thermal conductivity. Sogabe et al. (?) developed a small soft actuator that operates in environments rich in thermal energy, such as warm water. This innovation improved the actuation time for bending to around one second using only environmental energy. However, this actuator leverages the thermal properties of the surrounding water (specific heat capacity of 4186 $\rm{J/kg\cdot K}$ (?)) and is thus limited to underwater use.
To enable use in more diverse environments, systems capable of operating on land are required. A solar-driven actuator presents a promising approach to solving these issues  (?). Shao et al. proposed an actuator that absorbs environmental light energy through blackbody radiation by mixing graphite into silicone, converting it into thermal energy, and driving it using the liquid-to-gas phase change (methanol) (?). However, this method results in the loss of actuator transparency due to the use of graphite, and adding powders like graphite reduces the Young’s modulus, leading to constraints on the actuator’s shape and functionality.
Therefore, developing a mechanism that efficiently absorbs environmental energy, enabling actuation even with highly transparent silicone actuators, is crucial. We introduce mechanisms inspired by biological systems that efficiently absorb environmental energy into the actuator.
Bio-inspiration for efficient use of environmental energy
Nocturnal animals need to hunt and escape predators under the faint light of the moon and stars. To efficiently transmit this minimal light to the retinal photoreceptors, they possess a special structure called the tapetum lucidum (Fig. 1A). This structure captures light that passes through the photoreceptors and reflects it back to the retinal layer, stimulating the photoreceptors once again. Inspired by the relationship between the tapetum and the retina, which ‘recovers and reuses’ light that would otherwise escape, our study proposes the introduction of a new light-reusing bilayer structure that enables liquid-to-gas phase change, which is inefficient with silicone layers alone.
In this study, we employed to use Laser Induced Graphene (LIG) as the photothermal conversion element for the ”tapetum” layer. In recent years, Laser-Induced Graphene (LIG) has emerged as a prominent research focus in the field of direct laser writing techniques, where organic materials such as polyimide are transformed into graphene-like structures. LIG is characterized by the formation of a porous three-dimensional graphene network structure, produced through rapid thermal carbonization (pyrolysis) via laser irradiation. Lin et al. proposed a method to directly graphenize polyimide film surfaces using $\rm{CO_{2}}$  laser irradiation (?). Subsequently, Chyan et al. (?) demonstrated the versatility of LIG fabrication on diverse substrates including textiles, paper, and food products through repeated laser exposure, thereby accelerating its application in flexible devices. LIG has garnered significant attention as a photothermal conversion material that requires no electrical connections or external wiring. It exhibits efficient light-to-heat conversion properties under laser or lamp irradiation and demonstrates transferability to elastomeric substrates. For instance, Dallinger et al. developed an actuator by transferring LIG onto PDMS-based hydrogel sheets, where the generated heat induced humidity changes in the sheet, resulting in strain-induced bending (?). These characteristics have positioned LIG as a promising material for applications in bioelectronics and soft actuators (?).
Building upon these advances, we focused on expanding LIG’s applications to liquid-to-gas phase change actuators. We added a LIG layer to conventional silicone soft actuators (?) to enable efficient light-driven actuation. We propose a novel anisotropic bilayer structure where a LIG layer is positioned on the light-incident surface inside highly transparent silicone. This design, inspired by the eyes of nocturnal animals, enables efficient light collection and photothermal conversion. The converted thermal energy is transferred to the internal fluid, causing deformation of the silicone actuator and resulting in controlled bending motion. In this study, we present the fabrication process of the proposed actuator and evaluate its performance characteristics. Furthermore, to demonstrate the practical applicability of our system, we developed a multi-finger robot where multiple units of these novel actuators operate synchronously in response to light stimuli.
Figure 1: Concept and background of the proposed actuator. A: Structure of tapetum lucidum in nocturnal animals’ eyes and its application to the proposed actuator concept utilizing enhanced light absorption efficiency and photothermal conversion as driving force. Inset shows an actual photograph of feline tapetum lucidum (white arrow). B: Schematic illustration of the actuator’s driving mechanism and operational principles. C: Photograph of the fabricated actuator prototype. Scale bar = 10 mm. D: Comparative analysis between conventional soft actuators utilizing black materials (e.g., graphene) for photothermal conversion and the proposed actuator design, highlighting key differences.
2 Concept and fabrication process
2.1 Concept
Figure 1 illustrates the conceptual design of our proposed actuator. While previous phase-change actuators typically employed homogeneous structures, our novel approach focuses on the layered wall architecture to achieve efficient light absorption. As previously discussed, biological eyes utilize a structure called the tapetum to reflect escaped light, thereby increasing photoreceptor stimulation. Similarly, our design incorporates a Laser-induced graphene (LIG) layer to capture light not absorbed by the conventional silicone elastomer layer, converting it to thermal energy through photothermal conversion to efficiently vaporize and expand the internal low-boiling-point liquid (Fig. 1A).
The actuation mechanism of the proposed device is depicted in Fig. 1B. The soft actuator deforms through the vaporization of an internal low-boiling-point liquid, controlled by an external light source. The actuator is designed with asymmetric silicone wall thicknesses, with the thinner wall on the graphene layer side to induce directional bending toward the LIG layer. The operating temperature was set at 33C, above room temperature (25C), and we selected an inert fluorinated liquid (Opteon SF33, Chemours Company, base compound structure: $\rm{Cis-CF_{3}CH=CHCF_{3}}$) as the working fluid.
Figure 1C shows a photograph of the fabricated actuator. The actuator has an overall length of 20 mm with a flattened elliptical cross-section measuring approximately 5 mm along the major axis and 4 mm along the minor axis. As illustrated in Fig. 1D, our proposed actuator achieves a layered structure using LIG without incorporating thermally conductive materials or photothermal converters directly into the silicone walls. This approach not only facilitates efficient heat transfer but also maintains the transparency of the silicone elastomer, enabling broader applications of the elastomer-based system.
2.2 Examination of LIG layer characteristics
2.2.1 LIG Transfer Process onto Silicone
The LIG patterns are generated on Kapton tape (3 M Polyimide Film Tape 5413, 3 M Company, Minnesota, USA) adhered to a slide glass. The polyimide base of the Kapton tape enables LIG formation through laser irradiation. We utilize a laser system (xTool F1, Makeblock Co., Ltd., Guangdong, China) with specific parameters (200 mm/s speed, 40% power, unidirectional) for LIG patterning.
The LIG sheet was transferred onto silicone (Ecoflex 00-45 Near Clear, Smooth-On, Inc., Pennsylvania, USA). For the transfer process, the LIG sheet, which remained adhered to the slide glass on one side, was positioned and surrounded by a mold. The degassed uncured silicone was then poured into the mold using a vacuum degassing chamber. The curing process was carried out at 60C for 60 minutes. After curing, the mold was separated from the slide glass and the cured silicone was peeled off the Kapton tape, resulting in the successful transfer of the LIG onto the silicone surface, yielding a silicone substrate with an integrated LIG layer.
Figure 2: Fabrication process of the proposed actuator. In step (i), a laser-processed plate with LIG printing on slide glass is integrated with a mold component designed to form the tubular structure. Subsequently, in step (ii), a metallic rod (2.5 mm in diameter) is inserted and secured at the rod fixation port, followed by the injection of degassed Ecoflex 00-45. After curing at 60C for 60 minutes, step (iii) involves carefully removing the mold and peeling the silicone tube from the plate. At this stage, the LIG layer is positioned on the exterior surface; therefore, in step (iv), the tube is inverted. The fabrication process concludes with step (v), where one end of the open-ended tube is sealed with uncured silicone. The actuator becomes operational upon encapsulation with Opteon SF33 and securing the open end with a silicone-coated clip.
2.3 LIG-transferred soft actuator fabrication process
Figure 2 illustrates the fabrication process of a soft actuator with the LIG layer positioned adjacent to the liquid interface. The fabrication process consists primarily of five steps, designated as (i)–(v). Initially, a LIG pattern (2.6 mm $\times$ 40 mm) is created using the aforementioned laser parameters. The pattern is then positioned and secured within the mold (detailed in Supplementary Data S1) as shown in (i-a). Subsequently, a metallic rod (2.5 mm in diameter) is inserted into the mold, and silicone is introduced and cured as depicted in (ii). In step (iii), the tubular silicone structure with the integrated LIG layer is carefully removed from the mold. Because the LIG layer is initially present on the external surface, the tube is inverted in step (iv) to position the LIG layer on the interior. In step (v), uncured silicone is introduced into one end of the tube to create a seal, resulting in an actuator with one open terminus. The actuator is completed by introducing 40 µL of a low-boiling-point liquid (Opteon SF33, The Chemours Company, Delaware, USA) into the internal cavity and sealing the open end with a silicone-coated clip.
Additionally, the silicone sheets (10 mm square with transferred LIG patterns) were fabricated using the same process with molds adjusted to achieve a thickness of 1.0 mm. These molds were fabricated using a Formlabs Form 3 printer (Formlabs Inc., Somerville, MA, USA). Clear Resin v4 (Formlabs Inc.) was utilized as the resin material for mold production.
To investigate whether placing the silicone layer on the exterior and the LIG layer on the interior would affect LIG’s photothermal conversion efficiency, we examined the temperature response characteristics. A reflector lamp (RF100 V54 WD, 100 V, 54 W; Panasonic Corporation, Osaka, Japan) was used as the light source, and temperature changes on the silicone sheet were measured using a thermal camera (OPTPI64 ILTO15 T090, Optris GmbH, Berlin, Germany).
Our findings revealed that positioning LIG on the interior actually accelerated the temperature rise at the liquid-contact interface (Supplementary Fig. 1). The initial temperature change rate during the first 5 seconds after light irradiation was 0.71 K/s when irradiated from the silicone side, compared to 0.66 K/s when irradiated from the LIG side (reference value: approximately 0.13 K/s for silicone-only sheet). This phenomenon can be attributed to silicone’s high volumetric heat capacity and low thermal conductivity, which impede the efficient transfer of heat generated by LIG to the liquid-contact interface.
Figure 3: Durability assessment of LIG-transferred silicone under various mechanical and chemical stresses. (A) Resistance to repetitive stretching deformation, (B) resistance to abrasive wear, and (C) resistance to liquid exposure. No delamination or degradation of LIG was observed in any of the test conditions. Scale bar = 10 mm.
2.4 Mechanical and chemical durability of LIG-transferred silicone
Figure 3 illustrates the silicone sheet (10 mm $\times$ 10 mm) with a transferred LIG layer, as well as the condition of the LIG after being subjected to 20 cycles of stretching and relaxation, and twisting, respectively. In both cases, no delamination of the LIG layer was observed.
The resistance of the actuator to the liquid was evaluated by immersing the LIG-transferred sheet into a conical tube filled with Opteon SF33. The actuator was agitated for 60 seconds. After the process, the sheet was examined and no detachment or degradation of the LIG layer was observed, confirming its stability under these conditions (Fig. 4).
3 Heat transfer analysis
When an object is heated by radiation from a heat source at temperature $\theta_{h}$, the radiative heat transfer $Q_{hs}$ to the silicone wall can be expressed as
$Q_{hs}=\frac{\sigma(\theta_{h}^{4}-\theta_{s}^{4})\,A}{\dfrac{1}{\epsilon_{h}}+\dfrac{1}{\epsilon_{s}}-1}.$ (1)
Here, $\sigma=5.67037442\times 10^{-8}~{}\mathrm{W\cdot m^{-2}\cdot K^{-4}}$ is the Stefan-Boltzmann constant; $\theta_{s}$ is the temperature of the silicone wall; $A$ is the heat transfer area; and $\epsilon_{h}$ and $\epsilon_{s}$ are the emissivities of the heat source and the silicone wall, respectively ( (?)). The thermal conductivity of the silicone, $\lambda_{s}$, and its thickness, $x$, are given in Table 1  (?).
The actuator’s response time is sufficiently long. Therefore, the thermal Fourier number becomes large ($F_{o}>>1$) and the temperature distribution within the silicone wall is negligible. The energy balance equation for the silicone wall can then be written as
[…]
where $C_{s}$ is the specific heat capacity of the silicone, $W_{s}$ is its mass ($W_{s}=\rho_{s}\,A\,x$), $h_{se}$ is the heat transfer coefficient between the silicone wall and the environment, and $\theta_{e}$ is the environmental temperature (25C).
Using a difference method, the following equation is derived from Eqs. (1) and (2):
$\begin{split}\theta_{s}(t+\Delta t)&=\theta_{s}(t)+\frac{\sigma\,[\,\theta_{h}^{4}-\theta_{s}(t)^{4}\,]\;A}{C_{s}\,\rho_{s}\,A\,x\,\Bigl{(}\dfrac{1}{\epsilon_{h}}+\dfrac{1}{\epsilon_{s}}-1\Bigr{)}}-\frac{2\,h_{se}\,\bigl{[}\theta_{s}(t)-\theta_{e}\bigr{]}}{C_{s}\,\rho_{s}\,x}.\end{split}$ (3)
When the LIG layer is deposited on the side of the silicone wall opposite the heat source, the heat transfer from the heat source to the LIG layer is
$Q_{hL}=\frac{\sigma\,(\theta_{h}^{4}-\theta_{L}^{4})\,A}{\dfrac{1}{\epsilon_{h}}+\dfrac{1}{\epsilon_{L}}-1},$ (4)
where $\theta_{L}$ is the temperature of the LIG layer and $\epsilon_{L}$ is its emissivity. The energy balance equation for the LIG layer is
$C_{L}W_{L}\frac{d\theta_{L}}{dt}=Q_{hL}\;-\;h_{Le}\,A\,(\theta_{L}-\theta_{e})\;-\;Q_{Ls},$ (5)
with
$Q_{Ls}\simeq\lambda_{s}\,A\,\frac{\theta_{L}-\theta_{s}}{x}.$ (6)
Here, $C_{L}$ is the specific heat capacity of the LIG layer, $W_{L}=\rho_{L}\,A\,x_{L}$ is its mass, $h_{Le}$ is the heat transfer coefficient between the LIG layer and the environment, and $\lambda_{s}$ is the thermal conductivity of the silicone. Under these conditions, the energy equation of the silicone wall is written as
$C_{s}W_{s}\frac{d\theta_{s}}{dt}=Q_{hs}\;-\;\,h_{se}\,A\,(\theta_{s}-\theta_{e})\;+\;Q_{Ls}.$ (7)
The temperature of the LIG layer, $\theta_{L}(t)$, can be computed by a difference method based on Eqs. (4)–(7).
The parameters used in the simulation are summarized in Table 1 (?). (?)
Table 1: Parameters used in the simulation
Parameter Value
$\theta_{s}(0),\theta_{L}(0),\theta_{e}$ $298~{}\rm{K}$
$h_{se}$ $6~{}\rm{W/m^{2}\cdot K}$
$h_{Le}$ $25~{}\rm{W/m^{2}\cdot K}$
$C_{s}$ $1300~{}\rm{J/kg\cdot K}$
$C_{L}$ $700~{}\rm{J/kg\cdot K}$
$\rho_{s}$ $1050~{}\rm{kg/m^{3}}$  (?)
$\rho_{L}$ $400~{}\rm{kg/m^{3}}$
$\lambda_{s}$ $0.2~{}\rm{W/m\cdot K}$
$\alpha_{s}$ $0.17$
$\alpha_{L}$ $0.83$
$Q_{h}$ $0.1~{}\rm{W}$
In this simulation, for the purpose of model simplification, several factors such as Opteon SF33 vaporization and silicone elastomer elongation were not taken into account. The numerical values used for verification were obtained from the referenced literature. Additionally, while the radiative heat transfer equations are presented above, in our actual simulation we simplified the heat transfer calculation by considering the heat source as a light source with constant energy $Q_{h}$. The energy absorbed by each material was then calculated using material-specific absorption coefficients($\alpha_{s}$ and $\alpha_{L}$) which are determined experimentally, where $\alpha_{s}Q_{h}$ represents the energy absorbed by the silicone wall and $\alpha_{L}Q_{h}$ represents the energy absorbed by the LIG layer.
For comparison with simulation results, we measured the temperature at the liquid contact surface using a thermal camera with silicone sheets of the same thickness as in the simulation (The experimental setup for thermal data acquisition is shown in Fig. 4A, while the corresponding thermal imaging results are presented in Fig. 4B). The results are shown in Fig. 4C. When overlaying the temporal temperature variations with the simulation graphs, despite some discrepancies, the T63 values showed close agreement: for the case without LIG, the simulation yielded 118.1 seconds while the experimental measurement was 110.7 seconds; for the case with LIG, the simulation predicted 53.9 seconds compared to the experimental value of 58.3 seconds. These results demonstrate that LIG effectively functions as a photothermal conversion and facilitates heat transfer to the internal liquid.
Figure 4: Thermodynamic analysis of temperature changes at liquid-solid interface through thermal imaging and simulation A: Experimental setup for thermal imaging analysis.The measurement was conducted with the distance between the light source and the silicone sheet set to 100 mm. B: Real-time thermal imaging of temperature elevation at the liquid-solid interface comparing silicone sheets with- and without-LIG. The LIG-transferred silicone sheet demonstrated more rapid temperature elevation and higher maximum temperature (334.9 K with-LIG vs. 311.5 K without-LIG). C: Comparison between experimental thermal imaging data (solid lines) and thermodynamic simulation results (dashed lines). Time constant (T63) analysis revealed simulated T63 values of 118.1 sec and 53.9 sec for without- and with-LIG groups, respectively, while experimental measurements yielded T63 values of 110.7 sec and 58.3 sec, respectively.
Figure 5: Overview of the proposed actuator and experimental setup. A: Schematic illustration and dimensions of the proposed actuator. The cross-sectional view reveals a distorted elliptical shape, with slight deformation observed at the adhesion interface with the LIG. B: Schematic diagram of the light irradiation experimental setup. The fundamental experiments were conducted with the actuator positioned at a distance of 50 mm from the light source. A 54 W reflector bulb operating at 100 V served as the light source. C: Measurement method of bending angle: The initial state of the actuator before bending was defined as 0 degrees, and the angle by which the tip of the actuator tilted from this position was recorded as the bending angle.
Figure 6: Temporal analysis of bending actuator dynamics. A: Time-lapse photographs illustrating bending deformation during light irradiation with and without LIG. B: Graph of the relative bending angle ratio during light irradiation (n=5). C: Time-lapse photographs depicting the bending recovery process following light deactivation, and D: Corresponding graphical analysis of the relative bending angle ratio changes (n=5). The relative bending angle ratio is defined as the relative angle when the actuator response reaches a plateau (defined as an increase of less than 1 degree over a 20-second proximity period) or the angle in the frame immediately preceding angle reduction, with this value normalized to 1.
4 Evaluation of a soft actuator with transferred LIG under general light source
4.1 Experimental analysis of actuation driven by light On-Off cycling
The experimental setup is illustrated in Fig. 5. The fabricated actuator was positioned 50 mm away from the light source during the experiments. The light bulb specifications include Central luminance is 300  cd, total luminous flux is 630 lm, and beam angle is 60°. An iPhone 15 Pro Max (Apple Inc., Cupertino, CA, USA) was used to capture the actuator deformation, and ImageJ (citation) was employed for analyzing the bending angles and deformation measurements.
Figure 6A and C shows the bending behavior during light irradiation and the subsequent recovery upon light off, along with the corresponding analysis results. When exposed to light, the without-LIG actuator exhibited delayed activation, with a 63% response time of 141.7 seconds and approximately 220 seconds required to reach maximum bending angle. In contrast, the proposed actuator with LIG demonstrated a mean 63% response time of 65.0 seconds and reached maximum bending angle in approximately 90 seconds, showing an 54.1% improvement in responsiveness compared to the without-LIG actuator.
Regarding the recovery phase after light off (Fig 6B and D, the without-LIG actuator showed a 63% response time of 42.7 seconds and required 120 seconds to return to its initial state. The proposed actuator with LIG demonstrated enhanced recovery characteristics with a 63% response time of 22.1 seconds and approximately 60 seconds to return to its initial state, representing an 48.2% improvement in recovery performance compared to the without-LIG actuator.
Figure 7: Analysis of LIG-transferred soft actuators under different internal and environmental conditions: A-C: When the distance from the light source was constant (50 mm), the deformation ratio (A), plateau angle (B), and time to reach plateau state (C) were measured for internal liquid volumes of 20, 30, and 40 µL. D-F: When the internal liquid volume was constant (30 µL), the deformation ratio (D), plateau angle (E), and time to reach plateau state (F) were measured for distances of 50, 75, and 100 mm between the light source and actuator. The plateau state was defined as the condition where the angular variation did not increase by more than 1° over a 30-second period. The verification was conducted using three actuators (n=3) that exhibited similar plateau angles when operated with an internal volume of 20 µL at a distance of 50 mm from the light source.
4.2 Deformation of the actuator before and after light-driven operation
Figure 7 demonstrates the variations in actuator performance based on changes in the encapsulated liquid volume and the distance between the light source and actuator. With the light source fixed at 50 mm, we examined the effect of varying volumes (20 µL, 30 µL, and 40  µL) of encapsulated Opteon SF33. The results showed volume-dependent increases in three parameters: the rate of actuator deformation (Fig. 7A), the plateau bending angle (Fig. 7B), and the time required to reach the plateau state (Fig. 7C). The linear regression analysis yielded an R² value of approximately 0.98, indicating that the relationship can be effectively approximated as linear within the 20-40 µL range.
The influence of external environmental conditions was investigated by varying the distance between the light source and the actuator (50 mm, 75 mm, and 10 mm), with results presented in Fig. 7D-F (Opteon SF33’s amout is 30 µL). The deformation ratio and plateau bending angle decreased as the distance increased. Notably, the time required to reach the plateau state was longest in the 100 mm group, showing more than a 2.5-fold increase compared to the 50 mm group.
Figure 8: Demonstration of a bio-inspired robot utilizing the proposed actuator in a biomimetic movement driven by light. A: Venus flytrap used as the biomimetic target. B: Conceptual demonstration of the movement, where the robot curves to enclose and retain an object beneath it under light irradiation, and releases it when the light is turning off. C: Actual operation of the robot in the environment, using the same light source as in the experiments. Upon light exposure, the robot captures a miniature ladybug model, protecting and holding it from the light. When irradiation ceases, the robot releases its curvature and frees the enclosed object.
5 Multi-Finger Robot Using Non-Contact & Light-Driven LIG-Transferred Soft Actuators
To verify the synchronous actuation capability of multiple LIG-transferred soft actuators, we developed a three-fingered robot inspired by the Venus flytrap and evaluated its operation in ambient conditions. The demonstration is shown in Movie S2 and Fig. 8. As illustrated in Fi. 8, upon irradiation from a light source (a lamp in this experiment), the actuators began to bend synchronously, successfully enveloping a ladybug-inspired target object through their coordinated curving motion. When the light source was switched off, each actuator released from its bent state and returned to its original configuration.
6 Limitations & future work
6.1 The operating temperature range control
For the actuator’s internal fluid, we employed Opteon SF33 with a boiling point of 33C. We hypothesize that the operating temperature range can be controlled by varying the boiling point of the encapsulated liquid. For instance, Opteon SF79 and Opteon SF70 have boiling points of 48C and 71C, respectively. Thermal imaging data (Fig. 4B) demonstrates that temperatures approaching 55C can be achieved when LIG is transferred, suggesting that for terrestrial applications, this system could be implemented with liquids having boiling points in the range of 30-55C. However, for operations requiring higher temperature ranges, conventional light sources may prove insufficient. In such cases, the implementation of focusing lenses, high-powered lighting systems, or higher-output lasers is required.
6.2 Gas leakage
As shown in Fig. 7F, increasing the distance from the light source resulted in a decrease in the bending angle at plateau. This suggests that Opteon SF33 is leaking out due to silicone’s high gas permeability during the slower actuation caused by reduced light intensity. Similar gas leakage phenomena have been discussed in previous studies (?). The current actuator design implements a reusable mechanism where one side is sealed with a clip, allowing multiple refilling of the liquid. However, frequent leakage can limit potential applications. To address this limitation, potential approaches include the incorporation of materials with lower gas permeability and the development of alternative sealing mechanisms.
6.3 Utilization of ambient environmental light
The light source utilized in this study was a reflector lamp (54W) with a beam angle of 60°, center luminous flux of 630 lm, and beam flux of 160 lm. Assuming uniform light distribution, calculations based on illuminance indicate that the actuator surface, positioned at a distance of 50 mm, receives approximately 60,000 lux (assuming an actuator LIG surface area of 40 $\rm{mm}^{2}$). This illuminance corresponds to approximately 60% of direct summer sunlight (?). While outdoor operation would be dependent on seasonal and geographical factors, the implementation of lens-based light concentration mechanisms could make outdoor applications feasible. Given that typical office environments maintain illuminance levels of approximately 300–1000 lux (?), indoor operation under ambient lighting conditions would be significantly limited. Future investigations will explore localized light energy irradiation using laser sources, aimed at achieving more precise actuator control at a microscale level.
7 Conclusions
In this paper, we focused on the ocular structure of nocturnal animals that efficiently collect and utilize light in low-irradiation environments. We addressed the limitation of conventional silicone-based soft actuators, which exhibited low driving efficiency due to ineffective light utilization, by incorporating a light-heat conversion element layer called LIG on the light-incident surface. This modification enabled the actuator to effectively utilize light that would otherwise be poorly absorbed by silicone alone, thereby heating the internal low-boiling-point liquid. As a result, the proposed actuator with LIG demonstrated a 54% improvement in response time compared to conventional without-LIG actuators.
While various approaches have successfully enhanced photothermal conversion and thermal conductivity through the incorporation of graphene or metal powders into silicone elastomers, our method explores an alternative approach by installing a LIG layer on the silicone surface. This approach preserves the inherent properties of silicone while utilizing LIG, which can be manufactured under standard environmental conditions, thereby facilitating both the production of photothermal actuators and their diverse applications.
Future research will focus on the versatility of LIG transfer patterns, aiming to develop efficient light-utilizing LIG patterns under natural environmental conditions and design more complex actuator geometries with corresponding LIG patterns. This direction of research aligns with the principles of soft robotics, pursuing the development of soft actuators with enhanced deformation capabilities and environmental adaptability.
Acknowledgement
This work was supported by JSPS KAKENHI Grant Number JP24K21213. I sincerely thank my cats, whose eyes served as a visual reference for the demonstration of the tapetum lucidum structure and whose presence brought inspiration and support throughout the research process.
Author declaration
No competing financial interests exist.
CRediT
Maina Sogabe: Conceptualization, Data curation, Investigation, Formal analysis, Methodology,Supervision, Visualization, Writing – original draft, Writing – review & editing
Youhyun Kim: Data curation, Methodology, Formal analysis, Resources, Writing – original draft, Writing – review & editing
Kenji Kawahima: Funding acquisition, Investigation,Formal analysis,Project administration, Supervision, Writing – original draft, Writing – review & editing
8 Supplementary Material
  1. 1.
    Supplementary Figure S1:Changes in temperature icrease at liquid-interface surfaces with directional light irradiation: Thermographic analysis of temperature increase at liquid-contact interfaces comparing two irradiation patterns on LIG-transferred silicone sheets: (A) irradiation from silicone side to LIG side (proposed method) and (B) irradiation from LIG side to silicone side , with (C) showing reference values for silicone-only sheets. The proposed method with irradiation from the silicone side demonstrated higher initial temperature change rate (first 5 seconds after light exposure).
  2. 2.
    Supplementary Movie S1 :The actuator’s bending under light irradiation and its bending release upon turning off the light.
  3. 3.
    Supplementary Movie S2 :Demonstration of a bio-inspired (Venus flytrap) robot utilizing the proposed actuator in a biomimetic movement driven by light.

### v2

@@
-Nocturnal Eye–Inspired Liquid–to-Gas Phase Change Soft Actuator with Laser-Induced-Graphene : Enhanced Environmental Light Harvesting and Photothermal Conversion
+Nocturnal Eye-Inspired Liquid-to-Gas Phase Change Soft Actuator with Laser-Induced Graphene: Enhanced Environmental Light Harvesting and Photothermal Conversion
 Abstract
-The mobility of robotic systems is fundamentally constrained by their power sources and wiring requirements.
+Robotic systems’ mobility is fundamentally constrained by their power sources and wiring requirements.
 While electrical actuation systems have achieved autonomy through battery power and wireless control, pneumatic actuators remain tethered to air supply sources.
@@
 External heat sources, particularly light energy, present an alternative for terrestrial applications.
-However, silicone-based materials, despite their optical transparency, have high volumetric heat capacity and low thermal conductivity, limiting efficient photothermal energy transfer.
+However, despite their optical transparency, silicone-based materials have high volumetric heat capacity and low thermal conductivity, limiting efficient photothermal energy transfer.
 Previous attempts to address this through graphene or metallic powder incorporation have compromised material properties, including reduced transparency and altered elastic moduli.
@@
 This creates an anisotropic structure with enhanced photothermal conversion capabilities while maintaining the advantageous properties of silicone.
-Comparative analysis demonstrates that the proposed actuator exhibits significantly higher photo-induced bending efficiency compared to conventional silicone-based actuators.
+Comparative analysis demonstrates that the proposed actuator exhibits significantly higher photo-induced bending efficiency than conventional silicone-based actuators.
 The response time (T63) improved by 54%, decreasing from 142 seconds for pure silicone to 65 seconds, with recovery response time showing a 48% improvement.
-This design maintains silicone’s transparency and flexibility while utilizing LIG, which can be fabricated under ambient conditions, facilitating both manufacturing and diverse applications.
+This design maintains silicone’s transparency and flexibility while utilizing LIG, which can be fabricated under ambient conditions, facilitating manufacturing and diverse applications.
+keywords:
+Laser-induced graphene, liquid-to-gas phase change actuator and photothermal conversion
 1 Introduction
 Currently, most robotic actuators rely on electrical energy.
-For example, although pneumatic actuators use air as the driving source, they rely on electrical energy for control and to supply pressed air  (?) (?).
+For example, although pneumatic actuators use air as the driving source, they rely on electrical energy to control and supply pressed air [1][2].
 Therefore, they cannot be considered completely independent of electrical energy.
@@
 Aiming to use energy sources other than electrical energy to drive soft robots can significantly contribute to diversifying actuation methods in robotics.
-Several methods have been proposed on using heating methods, such as laser irradiation  (?) (?), micorwave  (?) , heatsink  (?) and electromagnetic  (?) (?).
+Several methods have been proposed for using heating methods, such as laser irradiation [3][4], microwave [5], heatsink [6] and electromagnetic [7][8].
 Kim et al.
-developed an electrothermal pneumatic actuator using nanofiber mats, achieving a lightweight soft actuator that can operate without external devices by utilizing liquid evaporation  (?).
+developed an electrothermal pneumatic actuator using nanofiber mats, achieving a lightweight, soft actuator that can operate without external devices by utilizing liquid evaporation [9].
 However, this approach faces issues such as high power dependency and slow response times.
-Various proposals have been made, starting with actuators like pouch motors ( (?))( (?))( (?))( (?))that utilize liquid-to-gas phase change, including pneumatic actuators driven by ambient temperature ( (?)).
+Various proposals have been made, starting with actuators like pouch motors ([10])([11])([12])([13])that utilize liquid-to-gas phase change, including pneumatic actuators driven by ambient temperature ([14]).
 However, these actuators operate relatively slowly, with many functioning at frequencies below 0.1Hz.
-This is due to significant losses in converting environmental energy into driving energy and issues related to heat transfer and thermal conductivity.
-Sogabe et al.
-(?) developed a small soft actuator that operates in environments rich in thermal energy, such as warm water.
+This is due to significant losses in converting environmental energy into driving energy, heat transfer, and thermal conductivity issues.
+Sogabe et al.[15] developed a small soft actuator that operates in environments rich in thermal energy, such as warm water.
 This innovation improved the actuation time for bending to around one second using only environmental energy.
-However, this actuator leverages the thermal properties of the surrounding water (specific heat capacity of 4186 $\rm{J/kg\cdot K}$ (?)) and is thus limited to underwater use.
-To enable use in more diverse environments, systems capable of operating on land are required.
-A solar-driven actuator presents a promising approach to solving these issues  (?).
+However, this actuator leverages the thermal properties of the surrounding water (specific heat capacity of 4186 $\rm{J/kg\cdot K}$[16]) and is thus limited to underwater use.
+For broader environmental applications.., systems capable of operating on land are required.
+A solar-driven actuator presents a promising approach to solving these issues [17].
 Shao et al.
-proposed an actuator that absorbs environmental light energy through blackbody radiation by mixing graphite into silicone, converting it into thermal energy, and driving it using the liquid-to-gas phase change (methanol) (?).
-However, this method results in the loss of actuator transparency due to the use of graphite, and adding powders like graphite reduces the Young’s modulus, leading to constraints on the actuator’s shape and functionality.
-Therefore, developing a mechanism that efficiently absorbs environmental energy, enabling actuation even with highly transparent silicone actuators, is crucial.
+proposed an actuator that absorbs environmental light energy through blackbody radiation by mixing graphite into silicone, converting it into thermal energy, and driving it using the liquid-to-gas phase change (methanol)[18].
+However, this method results in the loss of actuator transparency due to the use of graphite.
+Adding powders like graphite reduces Young’s modulus, leading to constraints on the actuator’s shape and functionality.
+Therefore, it is crucial to develop a mechanism that efficiently absorbs environmental energy, enabling actuation even with highly transparent silicone actuators.
 We introduce mechanisms inspired by biological systems that efficiently absorb environmental energy into the actuator.
 Bio-inspiration for efficient use of environmental energy
-Nocturnal animals need to hunt and escape predators under the faint light of the moon and stars.
-To efficiently transmit this minimal light to the retinal photoreceptors, they possess a special structure called the tapetum lucidum (Fig.
+Nocturnal animals need to hunt and escape predators under the moon’s and stars’ faint light.
+To efficiently transmit this minimal light to the retinal photoreceptors, they possess a unique structure called the tapetum lucidum (Fig.
 1A).
-This structure captures light that passes through the photoreceptors and reflects it back to the retinal layer, stimulating the photoreceptors once again.
-Inspired by the relationship between the tapetum and the retina, which ‘recovers and reuses’ light that would otherwise escape, our study proposes the introduction of a new light-reusing bilayer structure that enables liquid-to-gas phase change, which is inefficient with silicone layers alone.
-In this study, we employed to use Laser Induced Graphene (LIG) as the photothermal conversion element for the ”tapetum” layer.
-In recent years, Laser-Induced Graphene (LIG) has emerged as a prominent research focus in the field of direct laser writing techniques, where organic materials such as polyimide are transformed into graphene-like structures.
+This structure captures light that passes through the photoreceptors and reflects it to the retinal layer, again stimulating the photoreceptors.
+Inspired by the relationship between the tapetum and the retina, which ’recovers and reuses light that would otherwise escape, our study proposes the introduction of a new light-reusing bilayer structure that enables liquid-to-gas phase change, which is inefficient with silicone layers alone.
+In this study, we employed Laser-Induced Graphene (LIG) as the photothermal conversion element for the "tapetum" layer.
+In recent years, Laser-Induced Graphene (LIG) has emerged as a prominent research focus in direct laser writing techniques, where organic materials such as polyimide are transformed into graphene-like structures.
 LIG is characterized by the formation of a porous three-dimensional graphene network structure, produced through rapid thermal carbonization (pyrolysis) via laser irradiation.
 Lin et al.
-proposed a method to directly graphenize polyimide film surfaces using $\rm{CO_{2}}$  laser irradiation (?).
-Subsequently, Chyan et al.
-(?) demonstrated the versatility of LIG fabrication on diverse substrates including textiles, paper, and food products through repeated laser exposure, thereby accelerating its application in flexible devices.
+proposed a method to directly graphenize polyimide film surfaces using $\rm{CO_{2}}$  laser irradiation[19].
+Subsequently, Chyan et al.[20] demonstrated the versatility of LIG fabrication on diverse substrates, including textiles, paper, and food products, through repeated laser exposure, thereby accelerating its application in flexible devices.
 LIG has garnered significant attention as a photothermal conversion material that requires no electrical connections or external wiring.
-It exhibits efficient light-to-heat conversion properties under laser or lamp irradiation and demonstrates transferability to elastomeric substrates.
+It exhibits efficient light-to-heat conversion properties under laser or lamp irradiation, demonstrating transferability to elastomeric substrates.
 For instance, Dallinger et al.
-developed an actuator by transferring LIG onto PDMS-based hydrogel sheets, where the generated heat induced humidity changes in the sheet, resulting in strain-induced bending (?).
-These characteristics have positioned LIG as a promising material for applications in bioelectronics and soft actuators (?).
+developed an actuator by transferring LIG onto PDMS-based hydrogel sheets, where the generated heat-induced humidity changes in the sheet, resulting in strain-induced bending[21].
+These characteristics have positioned LIG as a promising material for applications in bioelectronics and soft actuators[22].
 Building upon these advances, we focused on expanding LIG’s applications to liquid-to-gas phase change actuators.
-We added a LIG layer to conventional silicone soft actuators (?) to enable efficient light-driven actuation.
-We propose a novel anisotropic bilayer structure where a LIG layer is positioned on the light-incident surface inside highly transparent silicone.
+To enable efficient light-driven actuation, we added a LIG layer to conventional silicone soft actuators[15].
+We propose a novel anisotropic bilayer structure with a LIG layer on the light-incident surface inside highly transparent silicone.
 This design, inspired by the eyes of nocturnal animals, enables efficient light collection and photothermal conversion.
@@
 Furthermore, to demonstrate the practical applicability of our system, we developed a multi-finger robot where multiple units of these novel actuators operate synchronously in response to light stimuli.
-Figure 1: Concept and background of the proposed actuator.
-A: Structure of tapetum lucidum in nocturnal animals’ eyes and its application to the proposed actuator concept utilizing enhanced light absorption efficiency and photothermal conversion as driving force.
-Inset shows an actual photograph of feline tapetum lucidum (white arrow).
-B: Schematic illustration of the actuator’s driving mechanism and operational principles.
-C: Photograph of the fabricated actuator prototype.
-Scale bar = 10 mm.
-D: Comparative analysis between conventional soft actuators utilizing black materials (e.g., graphene) for photothermal conversion and the proposed actuator design, highlighting key differences.
 2 Concept and fabrication process
@@
 Figure 1 illustrates the conceptual design of our proposed actuator.
-While previous phase-change actuators typically employed homogeneous structures, our novel approach focuses on the layered wall architecture to achieve efficient light absorption.
-As previously discussed, biological eyes utilize a structure called the tapetum to reflect escaped light, thereby increasing photoreceptor stimulation.
+While previous phase-change actuators typically employed homogeneous structures, our novel approach focuses on the layered wall architecture for efficient light absorption.
+As previously discussed, biological eyes utilize the tapetum to reflect escaped light, thereby increasing photoreceptor stimulation.
 Similarly, our design incorporates a Laser-induced graphene (LIG) layer to capture light not absorbed by the conventional silicone elastomer layer, converting it to thermal energy through photothermal conversion to efficiently vaporize and expand the internal low-boiling-point liquid (Fig.
@@
 1B.
-The soft actuator deforms through the vaporization of an internal low-boiling-point liquid, controlled by an external light source.
+The soft actuator deforms by vaporizing an internal low-boiling-point liquid controlled by an external light source.
 The actuator is designed with asymmetric silicone wall thicknesses, with the thinner wall on the graphene layer side to induce directional bending toward the LIG layer.
@@
 Figure 1C shows a photograph of the fabricated actuator.
-The actuator has an overall length of 20 mm with a flattened elliptical cross-section measuring approximately 5 mm along the major axis and 4 mm along the minor axis.
+The actuator has an overall length of 20 mm with a flattened elliptical cross-section measuring approximately 5 mm along the central axis and 4 mm along the minor axis.
 As illustrated in Fig.
-1D, our proposed actuator achieves a layered structure using LIG without incorporating thermally conductive materials or photothermal converters directly into the silicone walls.
-This approach not only facilitates efficient heat transfer but also maintains the transparency of the silicone elastomer, enabling broader applications of the elastomer-based system.
+1D, our proposed actuator achieves a layered structure using LIG without directly incorporating thermally conductive materials or photothermal converters into the silicone walls.
+This approach facilitates efficient heat transfer and maintains the transparency of the silicone elastomer, enabling broader applications of the elastomer-based system.
 2.2 Examination of LIG layer characteristics
@@
 The polyimide base of the Kapton tape enables LIG formation through laser irradiation.
-We utilize a laser system (xTool F1, Makeblock Co., Ltd., Guangdong, China) with specific parameters (200 mm/s speed, 40% power, unidirectional) for LIG patterning.
-The LIG sheet was transferred onto silicone (Ecoflex 00-45 Near Clear, Smooth-On, Inc., Pennsylvania, USA).
+We utilize a laser system (xTool F1, Makeblock Co., Ltd., Guangdong, China) with specific parameters (200 mm/s speed, 40% power, unidirectional) for LIG patterning[23].
+The LIG sheet was transferred onto a silicone (Ecoflex 00-45 Near Clear, Smooth-On, Inc., Pennsylvania, USA).
 For the transfer process, the LIG sheet, which remained adhered to the slide glass on one side, was positioned and surrounded by a mold.
@@
 The curing process was carried out at 60C for 60 minutes.
-After curing, the mold was separated from the slide glass and the cured silicone was peeled off the Kapton tape, resulting in the successful transfer of the LIG onto the silicone surface, yielding a silicone substrate with an integrated LIG layer.
-Figure 2: Fabrication process of the proposed actuator.
-In step (i), a laser-processed plate with LIG printing on slide glass is integrated with a mold component designed to form the tubular structure.
-Subsequently, in step (ii), a metallic rod (2.5 mm in diameter) is inserted and secured at the rod fixation port, followed by the injection of degassed Ecoflex 00-45.
-After curing at 60C for 60 minutes, step (iii) involves carefully removing the mold and peeling the silicone tube from the plate.
-At this stage, the LIG layer is positioned on the exterior surface; therefore, in step (iv), the tube is inverted.
-The fabrication process concludes with step (v), where one end of the open-ended tube is sealed with uncured silicone.
-The actuator becomes operational upon encapsulation with Opteon SF33 and securing the open end with a silicone-coated clip.
+After curing, the mold was separated from the slide glass, and the cured silicone was peeled off the Kapton tape, successfully transferring the LIG onto the silicone surface and yielding a silicone substrate with an integrated LIG layer.
 2.3 LIG-transferred soft actuator fabrication process
 Figure 2 illustrates the fabrication process of a soft actuator with the LIG layer positioned adjacent to the liquid interface.
-The fabrication process consists primarily of five steps, designated as (i)–(v).
+The fabrication process consists primarily of five steps: (i)–(v).
 Initially, a LIG pattern (2.6 mm $\times$ 40 mm) is created using the aforementioned laser parameters.
 The pattern is then positioned and secured within the mold (detailed in Supplementary Data S1) as shown in (i-a).
-Subsequently, a metallic rod (2.5 mm in diameter) is inserted into the mold, and silicone is introduced and cured as depicted in (ii).
+Subsequently, a metallic rod (2.5 mm in diameter) is inserted into the mold, and silicone is introduced and cured, as depicted in (ii).
 In step (iii), the tubular silicone structure with the integrated LIG layer is carefully removed from the mold.
 Because the LIG layer is initially present on the external surface, the tube is inverted in step (iv) to position the LIG layer on the interior.
-In step (v), uncured silicone is introduced into one end of the tube to create a seal, resulting in an actuator with one open terminus.
+In step (v), uncured silicone is introduced into one tube end to create a seal, resulting in an actuator with one open terminus.
 The actuator is completed by introducing 40 µL of a low-boiling-point liquid (Opteon SF33, The Chemours Company, Delaware, USA) into the internal cavity and sealing the open end with a silicone-coated clip.
@@
 These molds were fabricated using a Formlabs Form 3 printer (Formlabs Inc., Somerville, MA, USA).
-Clear Resin v4 (Formlabs Inc.) was utilized as the resin material for mold production.
-To investigate whether placing the silicone layer on the exterior and the LIG layer on the interior would affect LIG’s photothermal conversion efficiency, we examined the temperature response characteristics.
+Clear Resin v4 (Formlabs Inc.) was used as the mold production resin material.
+We examined the temperature response characteristics to investigate whether placing the silicone layer on the exterior and the LIG layer on the interior would affect LIG’s photothermal conversion efficiency.
 A reflector lamp (RF100 V54 WD, 100 V, 54 W; Panasonic Corporation, Osaka, Japan) was used as the light source, and temperature changes on the silicone sheet were measured using a thermal camera (OPTPI64 ILTO15 T090, Optris GmbH, Berlin, Germany).
@@
 This phenomenon can be attributed to silicone’s high volumetric heat capacity and low thermal conductivity, which impede the efficient transfer of heat generated by LIG to the liquid-contact interface.
-Figure 3: Durability assessment of LIG-transferred silicone under various mechanical and chemical stresses.
-(A) Resistance to repetitive stretching deformation, (B) resistance to abrasive wear, and (C) resistance to liquid exposure.
-No delamination or degradation of LIG was observed in any of the test conditions.
-Scale bar = 10 mm.
 2.4 Mechanical and chemical durability of LIG-transferred silicone
-Figure 3 illustrates the silicone sheet (10 mm $\times$ 10 mm) with a transferred LIG layer, as well as the condition of the LIG after being subjected to 20 cycles of stretching and relaxation, and twisting, respectively.
+Figure 3 illustrates the silicone sheet (10 mm $\times$ 10 mm) with a transferred LIG layer, as well as the condition of the LIG after being subjected to 20 cycles of stretching and, relaxation, and twisting, respectively.
 In both cases, no delamination of the LIG layer was observed.
@@
 The actuator was agitated for 60 seconds.
-After the process, the sheet was examined and no detachment or degradation of the LIG layer was observed, confirming its stability under these conditions (Fig.
+After the process, the sheet was examined, and no detachment or degradation of the LIG layer was observed, confirming its stability under these conditions (Fig.
 4).
@@
 $Q_{hs}=\frac{\sigma(\theta_{h}^{4}-\theta_{s}^{4})\,A}{\dfrac{1}{\epsilon_{h}}+\dfrac{1}{\epsilon_{s}}-1}.$ (1)
-Here, $\sigma=5.67037442\times 10^{-8}~{}\mathrm{W\cdot m^{-2}\cdot K^{-4}}$ is the Stefan-Boltzmann constant; $\theta_{s}$ is the temperature of the silicone wall; $A$ is the heat transfer area; and $\epsilon_{h}$ and $\epsilon_{s}$ are the emissivities of the heat source and the silicone wall, respectively ( (?)).
-The thermal conductivity of the silicone, $\lambda_{s}$, and its thickness, $x$, are given in Table 1  (?).
+Here, $\sigma=5.67037442\times 10^{-8}~{}\mathrm{W\cdot m^{-2}\cdot K^{-4}}$ is the Stefan-Boltzmann constant; $\theta_{s}$ is the temperature of the silicone wall; $A$ is the heat transfer area; and $\epsilon_{h}$ and $\epsilon_{s}$ are the emissivities of the heat source and the silicone wall, respectively ([24]).
+The thermal conductivity of the silicone, $\lambda_{s}$, and its thickness, $x$, are given in Table 1 [25].
 The actuator’s response time is sufficiently long.
-Therefore, the thermal Fourier number becomes large ($F_{o}>>1$) and the temperature distribution within the silicone wall is negligible.
+Therefore, the thermal Fourier number becomes large ($F_{o}>>1$), and the temperature distribution within the silicone wall is negligible.
 The energy balance equation for the silicone wall can then be written as
@@
 The temperature of the LIG layer, $\theta_{L}(t)$, can be computed by a difference method based on Eqs. (4)–(7).
-The parameters used in the simulation are summarized in Table 1 (?). (?)
+The parameters used in the simulation are summarized in Table 1[26].[27]
 Table 1: Parameters used in the simulation
@@
               $C_{L}$               $700~{}\rm{J/kg\cdot K}$
-$\rho_{s}$ $1050~{}\rm{kg/m^{3}}$  (?)
+              $\rho_{s}$               $1050~{}\rm{kg/m^{3}}$[28]
               $\rho_{L}$               $400~{}\rm{kg/m^{3}}$
               $\lambda_{s}$               $0.2~{}\rm{W/m\cdot K}$
+              $x$               $1.0\times 10^{-3}~{}\rm{m}$
+              $x_{L}$               $1.0\times 10^{-4}~{}\rm{m}$
+              $A$               $1.0\times 10^{-4}~{}\rm{m^{2}}$
               $\alpha_{s}$               $0.17$
@@
               $Q_{h}$               $0.1~{}\rm{W}$
-In this simulation, for the purpose of model simplification, several factors such as Opteon SF33 vaporization and silicone elastomer elongation were not taken into account.
+In this simulation, for model simplification, several factors, such as Opteon SF33 vaporization and silicone elastomer elongation, were not taken into account.
 The numerical values used for verification were obtained from the referenced literature.
-Additionally, while the radiative heat transfer equations are presented above, in our actual simulation we simplified the heat transfer calculation by considering the heat source as a light source with constant energy $Q_{h}$.
-The energy absorbed by each material was then calculated using material-specific absorption coefficients($\alpha_{s}$ and $\alpha_{L}$) which are determined experimentally, where $\alpha_{s}Q_{h}$ represents the energy absorbed by the silicone wall and $\alpha_{L}Q_{h}$ represents the energy absorbed by the LIG layer.
+Additionally, while the radiative heat transfer equations are presented above, in our simulation, we simplified the heat transfer calculation by considering the heat source as a light source with constant energy $Q_{h}$.
+The energy absorbed by each material was then calculated using material-specific absorption coefficients($\alpha_{s}$ and $\alpha_{L}$), which are determined experimentally, where $\alpha_{s}Q_{h}$ represents the energy absorbed by the silicone wall and $\alpha_{L}Q_{h}$ represents the energy absorbed by the LIG layer.
 For comparison with simulation results, we measured the temperature at the liquid contact surface using a thermal camera with silicone sheets of the same thickness as in the simulation (The experimental setup for thermal data acquisition is shown in Fig.
@@
 These results demonstrate that LIG effectively functions as a photothermal conversion and facilitates heat transfer to the internal liquid.
-Figure 4: Thermodynamic analysis of temperature changes at liquid-solid interface through thermal imaging and simulation A: Experimental setup for thermal imaging analysis.The measurement was conducted with the distance between the light source and the silicone sheet set to 100 mm.
-B: Real-time thermal imaging of temperature elevation at the liquid-solid interface comparing silicone sheets with- and without-LIG.
-The LIG-transferred silicone sheet demonstrated more rapid temperature elevation and higher maximum temperature (334.9 K with-LIG vs.
-311.5 K without-LIG).
-C: Comparison between experimental thermal imaging data (solid lines) and thermodynamic simulation results (dashed lines).
-Time constant (T63) analysis revealed simulated T63 values of 118.1 sec and 53.9 sec for without- and with-LIG groups, respectively, while experimental measurements yielded T63 values of 110.7 sec and 58.3 sec, respectively.
-Figure 5: Overview of the proposed actuator and experimental setup.
-A: Schematic illustration and dimensions of the proposed actuator.
-The cross-sectional view reveals a distorted elliptical shape, with slight deformation observed at the adhesion interface with the LIG.
-B: Schematic diagram of the light irradiation experimental setup.
-The fundamental experiments were conducted with the actuator positioned at a distance of 50 mm from the light source.
-A 54 W reflector bulb operating at 100 V served as the light source.
-C: Measurement method of bending angle: The initial state of the actuator before bending was defined as 0 degrees, and the angle by which the tip of the actuator tilted from this position was recorded as the bending angle.
-Figure 6: Temporal analysis of bending actuator dynamics.
-A: Time-lapse photographs illustrating bending deformation during light irradiation with and without LIG.
-B: Graph of the relative bending angle ratio during light irradiation (n=5).
-C: Time-lapse photographs depicting the bending recovery process following light deactivation, and D: Corresponding graphical analysis of the relative bending angle ratio changes (n=5).
-The relative bending angle ratio is defined as the relative angle when the actuator response reaches a plateau (defined as an increase of less than 1 degree over a 20-second proximity period) or the angle in the frame immediately preceding angle reduction, with this value normalized to 1.
 4 Evaluation of a soft actuator with transferred LIG under general light source
@@
 5.
-The fabricated actuator was positioned 50 mm away from the light source during the experiments.
-The light bulb specifications include Central luminance is 300  cd, total luminous flux is 630 lm, and beam angle is 60°.
+During the experiments, the fabricated actuator was positioned 50 mm away from the light source.
+The light bulb specifications include luminous intensity of 300 cd, total luminous flux of 630 lm, and beam angle of $60^{\circ}$.
 An iPhone 15 Pro Max (Apple Inc., Cupertino, CA, USA) was used to capture the actuator deformation, and ImageJ (citation) was employed for analyzing the bending angles and deformation measurements.
-Figure 6A and C shows the bending behavior during light irradiation and the subsequent recovery upon light off, along with the corresponding analysis results.
-When exposed to light, the without-LIG actuator exhibited delayed activation, with a 63% response time of 141.7 seconds and approximately 220 seconds required to reach maximum bending angle.
-In contrast, the proposed actuator with LIG demonstrated a mean 63% response time of 65.0 seconds and reached maximum bending angle in approximately 90 seconds, showing an 54.1% improvement in responsiveness compared to the without-LIG actuator.
+Figure 6A and C show the bending behavior during light irradiation, the subsequent recovery upon light off, and the corresponding analysis results.
+When exposed to light, the without-LIG actuator exhibited delayed activation, with a 63% response time of 141.7 seconds and approximately 220 seconds required to reach the maximum bending angle.
+In contrast, the proposed actuator with LIG demonstrated a mean 63% response time of 65.0 seconds.
+It reached maximum bending angle in approximately 90 seconds, showing a 54.1% improvement in responsiveness compared to the without-LIG actuator.
 Regarding the recovery phase after light off (Fig 6B and D, the without-LIG actuator showed a 63% response time of 42.7 seconds and required 120 seconds to return to its initial state.
-The proposed actuator with LIG demonstrated enhanced recovery characteristics with a 63% response time of 22.1 seconds and approximately 60 seconds to return to its initial state, representing an 48.2% improvement in recovery performance compared to the without-LIG actuator.
-Figure 7: Analysis of LIG-transferred soft actuators under different internal and environmental conditions: A-C: When the distance from the light source was constant (50 mm), the deformation ratio (A), plateau angle (B), and time to reach plateau state (C) were measured for internal liquid volumes of 20, 30, and 40 µL.
-D-F: When the internal liquid volume was constant (30 µL), the deformation ratio (D), plateau angle (E), and time to reach plateau state (F) were measured for distances of 50, 75, and 100 mm between the light source and actuator.
-The plateau state was defined as the condition where the angular variation did not increase by more than 1° over a 30-second period.
-The verification was conducted using three actuators (n=3) that exhibited similar plateau angles when operated with an internal volume of 20 µL at a distance of 50 mm from the light source.
+The proposed actuator with LIG demonstrated enhanced recovery characteristics with a 63% response time of 22.1 seconds and approximately 60 seconds to return to its initial state, representing a 48.2% improvement in recovery performance compared to the without-LIG actuator.
 4.2 Deformation of the actuator before and after light-driven operation
@@
 The influence of external environmental conditions was investigated by varying the distance between the light source and the actuator (50 mm, 75 mm, and 10 mm), with results presented in Fig.
-7D-F (Opteon SF33’s amout is 30 µL).
+7D-F (Opteon SF33’s amount is 30 µL).
 The deformation ratio and plateau bending angle decreased as the distance increased.
-Notably, the time required to reach the plateau state was longest in the 100 mm group, showing more than a 2.5-fold increase compared to the 50 mm group.
-Figure 8: Demonstration of a bio-inspired robot utilizing the proposed actuator in a biomimetic movement driven by light.
-A: Venus flytrap used as the biomimetic target.
-B: Conceptual demonstration of the movement, where the robot curves to enclose and retain an object beneath it under light irradiation, and releases it when the light is turning off.
-C: Actual operation of the robot in the environment, using the same light source as in the experiments.
-Upon light exposure, the robot captures a miniature ladybug model, protecting and holding it from the light.
-When irradiation ceases, the robot releases its curvature and frees the enclosed object.
+Notably, the time required to reach the plateau state was the longest in the 100 mm group, showing more than a 2.5-fold increase compared to the 50 mm group.
 5 Multi-Finger Robot Using Non-Contact & Light-Driven LIG-Transferred Soft Actuators
@@
 8.
-As illustrated in Fi.
+As illustrated in Fig.
 8, upon irradiation from a light source (a lamp in this experiment), the actuators began to bend synchronously, successfully enveloping a ladybug-inspired target object through their coordinated curving motion.
-When the light source was switched off, each actuator released from its bent state and returned to its original configuration.
+When the light source was switched off, each actuator was released from its bent state and returned to its original configuration.
 6 Limitations & future work
 6.1 The operating temperature range control
-For the actuator’s internal fluid, we employed Opteon SF33 with a boiling point of 33C.
+We employed Opteon SF33 with a boiling point of 33C for the actuator’s internal fluid.
 We hypothesize that the operating temperature range can be controlled by varying the boiling point of the encapsulated liquid.
@@
 4B) demonstrates that temperatures approaching 55C can be achieved when LIG is transferred, suggesting that for terrestrial applications, this system could be implemented with liquids having boiling points in the range of 30-55C.
-However, for operations requiring higher temperature ranges, conventional light sources may prove insufficient.
-In such cases, the implementation of focusing lenses, high-powered lighting systems, or higher-output lasers is required.
+However, conventional light sources may prove insufficient for operations requiring higher temperature ranges.
+In such cases, focusing lenses, high-powered lighting systems, or higher-output lasers are required.
 6.2 Gas leakage
 As shown in Fig.
-7F, increasing the distance from the light source resulted in a decrease in the bending angle at plateau.
+7F, increasing the distance from the light source decreased the bending angle at a plateau.
 This suggests that Opteon SF33 is leaking out due to silicone’s high gas permeability during the slower actuation caused by reduced light intensity.
-Similar gas leakage phenomena have been discussed in previous studies (?).
-The current actuator design implements a reusable mechanism where one side is sealed with a clip, allowing multiple refilling of the liquid.
+Similar gas leakage phenomena have been discussed in previous studies[15].
+The current actuator design implements a reusable mechanism where one side is sealed with a clip, allowing multiple liquid refilling.
 However, frequent leakage can limit potential applications.
-To address this limitation, potential approaches include the incorporation of materials with lower gas permeability and the development of alternative sealing mechanisms.
+Potential approaches include incorporating materials with lower gas permeability and developing alternative sealing mechanisms to address this limitation.
 6.3 Utilization of ambient environmental light
-The light source utilized in this study was a reflector lamp (54W) with a beam angle of 60°, center luminous flux of 630 lm, and beam flux of 160 lm.
+The light source utilized in this study was a reflector lamp (54W) with a beam angle of $60^{\circ}$, center luminous flux of 630 lm, and beam flux of 160 lm.
 Assuming uniform light distribution, calculations based on illuminance indicate that the actuator surface, positioned at a distance of 50 mm, receives approximately 60,000 lux (assuming an actuator LIG surface area of 40 $\rm{mm}^{2}$).
-This illuminance corresponds to approximately 60% of direct summer sunlight (?).
-While outdoor operation would be dependent on seasonal and geographical factors, the implementation of lens-based light concentration mechanisms could make outdoor applications feasible.
-Given that typical office environments maintain illuminance levels of approximately 300–1000 lux (?), indoor operation under ambient lighting conditions would be significantly limited.
-Future investigations will explore localized light energy irradiation using laser sources, aimed at achieving more precise actuator control at a microscale level.
+This illuminance corresponds to approximately 60% of direct summer sunlight[29].
+While outdoor operation would depend on seasonal and geographical factors, implementing lens-based light concentration mechanisms could make outdoor applications feasible.
+Given that typical office environments maintain illuminance levels of approximately 300–1000 lux[30], indoor operation under ambient lighting conditions would be significantly limited.
+Future investigations will explore localized light energy irradiation using laser sources aimed at achieving more precise actuator control at a microscale level.
 7 Conclusions
-In this paper, we focused on the ocular structure of nocturnal animals that efficiently collect and utilize light in low-irradiation environments.
-We addressed the limitation of conventional silicone-based soft actuators, which exhibited low driving efficiency due to ineffective light utilization, by incorporating a light-heat conversion element layer called LIG on the light-incident surface.
-This modification enabled the actuator to effectively utilize light that would otherwise be poorly absorbed by silicone alone, thereby heating the internal low-boiling-point liquid.
+This paper focused on the ocular structure of nocturnal animals that efficiently collect and utilize light in low-irradiation environments.
+We addressed the limitation of conventional silicone-based soft actuators, which exhibited low driving efficiency due to ineffective light utilization.
+We incorporated a light-heat conversion element layer called LIG on the light-incident surface.
+This modification enabled the actuator to effectively utilize light that would otherwise be poorly absorbed by silicone alone, heating the internal low-boiling-point liquid.
 As a result, the proposed actuator with LIG demonstrated a 54% improvement in response time compared to conventional without-LIG actuators.
-While various approaches have successfully enhanced photothermal conversion and thermal conductivity through the incorporation of graphene or metal powders into silicone elastomers, our method explores an alternative approach by installing a LIG layer on the silicone surface.
+While various approaches have successfully enhanced photothermal conversion and thermal conductivity by incorporating graphene or metal powders into silicone elastomers, our method explores an alternative approach by installing a LIG layer on the silicone surface.
 This approach preserves the inherent properties of silicone while utilizing LIG, which can be manufactured under standard environmental conditions, thereby facilitating both the production of photothermal actuators and their diverse applications.
 Future research will focus on the versatility of LIG transfer patterns, aiming to develop efficient light-utilizing LIG patterns under natural environmental conditions and design more complex actuator geometries with corresponding LIG patterns.
-This direction of research aligns with the principles of soft robotics, pursuing the development of soft actuators with enhanced deformation capabilities and environmental adaptability.
+This research direction aligns with the principles of soft robotics, pursuing the development of soft actuators with enhanced deformation capabilities and environmental adaptability.
 Acknowledgement
 This work was supported by JSPS KAKENHI Grant Number JP24K21213.
-I sincerely thank my cats, whose eyes served as a visual reference for the demonstration of the tapetum lucidum structure and whose presence brought inspiration and support throughout the research process.
+M.S.
+sincerely thank my cats, whose eyes served as a visual reference for the demonstration of the tapetum lucidum structure and whose presence inspired and supported me throughout the research process.
 Author declaration
@@
 CRediT
-Maina Sogabe: Conceptualization, Data curation, Investigation, Formal analysis, Methodology,Supervision, Visualization, Writing – original draft, Writing – review & editing
-Youhyun Kim: Data curation, Methodology, Formal analysis, Resources, Writing – original draft, Writing – review & editing
-Kenji Kawahima: Funding acquisition, Investigation,Formal analysis,Project administration, Supervision, Writing – original draft, Writing – review & editing
+Maina Sogabe: Conceptualization, Data curation, Investigation, Formal analysis, Methodology, Supervision, Visualization, Writing-original draft, Writing-review & editing
+Youhyun Kim: Data curation, Methodology, Formal analysis, Resources, Writing-original draft, Writing-review & editing
+Hiroki Miyazako: Methodology, Writing-review & editing
+Kenji Kawashima: Funding acquisition, Investigation, Formal analysis, Project administration, Supervision, Writing-original draft, Writing-review & editing
 8 Supplementary Material
   1. 1.
-    Supplementary Figure S1:Changes in temperature icrease at liquid-interface surfaces with directional light irradiation: Thermographic analysis of temperature increase at liquid-contact interfaces comparing two irradiation patterns on LIG-transferred silicone sheets: (A) irradiation from silicone side to LIG side (proposed method) and (B) irradiation from LIG side to silicone side , with (C) showing reference values for silicone-only sheets.
-The proposed method with irradiation from the silicone side demonstrated higher initial temperature change rate (first 5 seconds after light exposure).
+    Supplementary Figure S1
   2. 2.
-    Supplementary Movie S1 :The actuator’s bending under light irradiation and its bending release upon turning off the light.
+    Supplementary Movie The actuator’s bending under light irradiation and its bending release upon turning off the light.
   3. 3.

### v4

@@
 While electrical actuation systems have achieved autonomy through battery power and wireless control, pneumatic actuators remain tethered to air supply sources.
-Liquid-to-gas phase change actuators utilizing low-boiling-point liquids offer a potential liquid, though they typically require substantial thermal input through heating elements that maintain electrical dependencies.
+Liquid-to-gas phase change actuators utilizing low-boiling-point liquids offer a potential solution, though they typically require substantial thermal input through heating elements that maintain electrical dependencies.
 External heat sources, particularly light energy, present an alternative for terrestrial applications.
-However, despite their optical transparency, silicone-based materials have high volumetric heat capacity and low thermal conductivity, limiting efficient photothermal energy transfer.
-Previous attempts to address this through graphene or metallic powder incorporation have compromised material properties, including reduced transparency and altered elastic moduli.
-Inspired by the tapetum lucidum structure found in nocturnal animals’ eyes for efficient light utilization in low-light conditions, this study proposes a novel anisotropic bilayer soft actuator incorporating Laser-Induced Graphene (LIG) on the inner surface of the light-irradiated silicone layer.
+However, despite their optical transparency, silicone-based materials have a high volumetric heat capacity and low thermal conductivity, which limits efficient photothermal energy transfer.
+Previous attempts to address this issue through the incorporation of graphene or metallic powder have compromised material properties, including reduced transparency and altered elastic moduli.
+Inspired by the tapetum lucidum structure found in the eyes of nocturnal animals, which enables efficient light utilization in low-light conditions, this study proposes a novel anisotropic bilayer soft actuator incorporating Laser-Induced Graphene (LIG) on the inner surface of the light-irradiated silicone layer.
 This creates an anisotropic structure with enhanced photothermal conversion capabilities while maintaining the advantageous properties of silicone.
 Comparative analysis demonstrates that the proposed actuator exhibits significantly higher photo-induced bending efficiency than conventional silicone-based actuators.
-The response time (T63) improved by 54%, decreasing from 142 seconds for pure silicone to 65 seconds, with recovery response time showing a 48% improvement.
-This design maintains silicone’s transparency and flexibility while utilizing LIG, which can be fabricated under ambient conditions, facilitating manufacturing and diverse applications.
+The response time improved by 54%, decreasing from 142 seconds for pure silicone to 65 seconds, with recovery response time showing a 48% improvement.
+This design maintains the silicone’s transparency and flexibility while utilizing LIG, which can be fabricated under ambient conditions, facilitating manufacturing and diverse applications.
 keywords:
@@
 Currently, most robotic actuators rely on electrical energy.
-For example, although pneumatic actuators use air as the driving source, they rely on electrical energy to control and supply pressed air [1][2].
+For example, although pneumatic actuators use air as the driving source, they rely on electrical energy to control and supply pressed air [1, 2].
 Therefore, they cannot be considered completely independent of electrical energy.
-Throughout history, humans have utilized various forms of energy to enrich their lives.
-Aiming to use energy sources other than electrical energy to drive soft robots can significantly contribute to diversifying actuation methods in robotics.
-Several methods have been proposed for using heating methods, such as laser irradiation [3][4], microwave [5], heatsink [6] and electromagnetic [7][8].
+For centuries, people have drawn upon a wide array of energy sources to enhance their lives.
+Exploring alternatives to electrical power for driving soft robots can greatly broaden the range of actuation techniques available in robotics.
+Several methods have been proposed for using heating methods, such as laser irradiation [3, 4], microwave [5], heatsink [6] and electromagnetic [7, 8].
 Kim et al.
@@
 However, this approach faces issues such as high power dependency and slow response times.
-Various proposals have been made, starting with actuators like pouch motors ([10])([11])([12])([13])that utilize liquid-to-gas phase change, including pneumatic actuators driven by ambient temperature ([14]).
+Various proposals have been made, starting with actuators like pouch motors [10, 11, 12, 13] that utilize liquid-to-gas phase change, including pneumatic actuators driven by ambient temperature [14].
 However, these actuators operate relatively slowly, with many functioning at frequencies below 0.1Hz.
@@
 However, this actuator leverages the thermal properties of the surrounding water (specific heat capacity of 4186 $\rm{J/kg\cdot K}$[16]) and is thus limited to underwater use.
-For broader environmental applications.., systems capable of operating on land are required.
-A solar-driven actuator presents a promising approach to solving these issues [17].
+For broader environmental applications, systems capable of operating on land are required.
+Thermal energy available in natural environments includes not only water temperature or body heat[17] but also ambient thermal gradients, which have been utilized for shape memory actuation and energy harvesting in soft robotic systems[18].
+In addition to these, other strategies harness photothermal conversion using sunlight as an energy source.
+A solar-driven approach utilizes a selective absorber film and a low-boiling-point fluid to generate internal pressure through sunlight-induced evaporation, enabling untethered soft robotic actuation [19].
 Shao et al.
 proposed an actuator that absorbs environmental light energy through blackbody radiation by mixing graphite into silicone, converting it into thermal energy, and driving it using the liquid-to-gas phase change (methanol)[20].
-However, this method results in the loss of actuator transparency due to the use of graphite.
-Adding powders like graphite reduces Young’s modulus, leading to constraints on the actuator’s shape and functionality.
-Therefore, it is crucial to develop a mechanism that efficiently absorbs environmental energy, enabling actuation even with highly transparent silicone actuators.
-We introduce mechanisms inspired by biological systems that efficiently absorb environmental energy into the actuator.
+However, using this approach leads to reduced actuator transparency because of the graphite content.
+When powders such as graphite are incorporated, they decrease Young’s modulus and impose limitations on both the actuator’s form and performance capabilities.
+For this reason, developing an efficient environmental energy absorption mechanism becomes essential - one that allows even highly transparent silicone actuators to function effectively.
+Our work presents bio-inspired mechanisms designed to efficiently capture environmental energy for actuator operation.
 Bio-inspiration for efficient use of environmental energy
-Nocturnal animals need to hunt and escape predators under the moon’s and stars’ faint light.
-To efficiently transmit this minimal light to the retinal photoreceptors, they possess a unique structure called the tapetum lucidum (Fig.
+Nocturnal animals rely on extremely low ambient light from the moon and stars to hunt and avoid predators.
+To maximize the capture of this scarce light, many nocturnal species possess a specialized reflective layer behind the retina called the tapetum lucidum (Fig.
 1A).
-This structure captures light that passes through the photoreceptors and reflects it to the retinal layer, again stimulating the photoreceptors.
-Inspired by the relationship between the tapetum and the retina, which ’recovers and reuses light that would otherwise escape, our study proposes the introduction of a new light-reusing bilayer structure that enables liquid-to-gas phase change, which is inefficient with silicone layers alone.
-In this study, we employed Laser-Induced Graphene (LIG) as the photothermal conversion element for the "tapetum" layer.
-In recent years, Laser-Induced Graphene (LIG) has emerged as a prominent research focus in direct laser writing techniques, where organic materials such as polyimide are transformed into graphene-like structures.
-LIG is characterized by the formation of a porous three-dimensional graphene network structure, produced through rapid thermal carbonization (pyrolysis) via laser irradiation.
-Lin et al.
-proposed a method to directly graphenize polyimide film surfaces using $\rm{CO_{2}}$  laser irradiation[19].
-Subsequently, Chyan et al.[20] demonstrated the versatility of LIG fabrication on diverse substrates, including textiles, paper, and food products, through repeated laser exposure, thereby accelerating its application in flexible devices.
-LIG has garnered significant attention as a photothermal conversion material that requires no electrical connections or external wiring.
-It exhibits efficient light-to-heat conversion properties under laser or lamp irradiation, demonstrating transferability to elastomeric substrates.
-For instance, Dallinger et al.
-developed an actuator by transferring LIG onto PDMS-based hydrogel sheets, where the generated heat-induced humidity changes in the sheet, resulting in strain-induced bending[21].
-These characteristics have positioned LIG as a promising material for applications in bioelectronics and soft actuators[22].
-Building upon these advances, we focused on expanding LIG’s applications to liquid-to-gas phase change actuators.
-To enable efficient light-driven actuation, we added a LIG layer to conventional silicone soft actuators[15].
-We propose a novel anisotropic bilayer structure with a LIG layer on the light-incident surface inside highly transparent silicone.
-This design, inspired by the eyes of nocturnal animals, enables efficient light collection and photothermal conversion.
-The converted thermal energy is transferred to the internal fluid, causing deformation of the silicone actuator and resulting in controlled bending motion.
-In this study, we present the fabrication process of the proposed actuator and evaluate its performance characteristics.
-Furthermore, to demonstrate the practical applicability of our system, we developed a multi-finger robot where multiple units of these novel actuators operate synchronously in response to light stimuli.
+By reflecting photons that have passed through the photoreceptors back into the retina, the tapetum lucidum significantly enhances photoreceptor stimulation under dim conditions.
+Conventional silicone-layer-only liquid-to-gas phase-change actuators absorb only a fraction of incident light, resulting in inefficient photothermal conversion.
+To improve light utilization efficiency, we use laser-induced graphene (LIG) to absorb light transmitted through the silicone layer and convert it into heat, analogous to how the tapetum lucidum recovers and reuses light that would otherwise be lost.
+LIG has become an important area of research in direct laser writing techniques.
+In this process, organic materials like polyimide are converted into porous, three-dimensional graphene-like structures through rapid thermal carbonization (pyrolysis) using $\rm{CO_{2}}$ laser irradiation [21, 22].
+The resulting porous network efficiently absorbs incident photons across a broad spectrum and converts them to thermal energy without requiring external wiring or electrical connections.
+Due to its high electrical conductivity and patternability, LIG has attracted considerable interest in flexible electronics and multifunctional sensing devices [23, 24].
+While most research has concentrated on electro-responsive or sensing applications, these studies reveal the expanding capabilities of LIG-based platforms.
+Wang et al.
+proposed light-driven robots utilizing the photothermal Marangoni effect [25].
+Furthermore, Dallinger et al.
+transferred LIG onto PDMS-based hydrogel sheets and achieved rapid, humidity-driven bending under light irradiation [26].
+These properties make LIG a promising candidate for bioelectronics and soft actuator applications [27].
+Building on these advances, we extended LIG’s application to liquid-to-gas phase-change actuators.
+To enable efficient light-driven actuation, we added a LIG layer to conventional silicone soft actuators [15], creating an anisotropic bilayer structure in which the LIG "tapetum" lies on the light-incident surface beneath highly transparent silicone.
+This design captures any light that passes through the silicone and converts it to thermal energy, which is then transferred to the internal low-boiling-point fluid, causing vaporization and controlled bending of the silicone actuator.
+In this study, we detail the fabrication process of the proposed actuator and evaluate its performance characteristics.
+Finally, to demonstrate practical applicability, we developed a multi-finger robot in which multiple LIG-enhanced actuators operate synchronously in response to light stimuli.
 2 Concept and fabrication process
@@
 Figure 1 illustrates the conceptual design of our proposed actuator.
-While previous phase-change actuators typically employed homogeneous structures, our novel approach focuses on the layered wall architecture for efficient light absorption.
-As previously discussed, biological eyes utilize the tapetum to reflect escaped light, thereby increasing photoreceptor stimulation.
-Similarly, our design incorporates a Laser-induced graphene (LIG) layer to capture light not absorbed by the conventional silicone elastomer layer, converting it to thermal energy through photothermal conversion to efficiently vaporize and expand the internal low-boiling-point liquid (Fig.
+Conventional silicone-layer-only liquid-to-gas phase-change actuators absorb only a fraction of incident light, leading to inefficient photothermal conversion.
+By analogy to the tapetum lucidum’s function of recovering and reusing otherwise lost light, our novel approach adopts a layered wall architecture to maximize light absorption.
+Specifically, we incorporate a Laser-Induced Graphene (LIG) layer on the light-incident surface of a highly transparent silicone elastomer (Fig.
 1A).
+This LIG "tapetum" captures any light that passes through the silicone, converting it into thermal energy via photothermal conversion to efficiently vaporize and expand the internal low-boiling-point liquid.
 The actuation mechanism of the proposed device is depicted in Fig.
 1B.
-The soft actuator deforms by vaporizing an internal low-boiling-point liquid controlled by an external light source.
-The actuator is designed with asymmetric silicone wall thicknesses, with the thinner wall on the graphene layer side to induce directional bending toward the LIG layer.
-The operating temperature was set at 33C, above room temperature (25C), and we selected an inert fluorinated liquid (Opteon SF33, Chemours Company, base compound structure: $\rm{Cis-CF_{3}CH=CHCF_{3}}$) as the working fluid.
-Figure 1C shows a photograph of the fabricated actuator.
-The actuator has an overall length of 20 mm with a flattened elliptical cross-section measuring approximately 5 mm along the central axis and 4 mm along the minor axis.
+Under external light irradiation, the LIG layer heats up and transfers thermal energy to the internal working fluid (Opteon SF33, Chemours Company; $\rm{Cis\text{-}CF_{3}CH=CHCF_{3}}$), causing it to vaporize.
+Because the silicone walls are purposely asymmetric—thinner on the LIG side—the actuator bends directionally toward the LIG layer when the fluid expands.
+The operating temperature is 33C (above room temperature, 25C), selected to match the boiling point of the chosen fluorinated liquid.
+Figure 1C shows a photograph of the fabricated actuator, which measures 20 mm in length with a flattened elliptical cross-section of approximately 5 mm by 4 mm.
 As illustrated in Fig.
-1D, our proposed actuator achieves a layered structure using LIG without directly incorporating thermally conductive materials or photothermal converters into the silicone walls.
-This approach facilitates efficient heat transfer and maintains the transparency of the silicone elastomer, enabling broader applications of the elastomer-based system.
+1D, our bilayer design achieves efficient heat transfer without embedding additional conductive or photothermal fillers into the silicone walls.
+The LIG layer is applied to only a small portion of the actuator’s interior, preserving transparency across most of the device’s area.
+This approach leaves most of the silicone structure clear while still taking advantage of LIG’s heat-generating properties when exposed to light.
+This design maintains the actuator’s lightweight and flexible nature while leveraging LIG’s photothermal properties and facilitating easy fabrication under ambient conditions.
 2.2 Examination of LIG layer characteristics
@@
 The polyimide base of the Kapton tape enables LIG formation through laser irradiation.
-We utilize a laser system (xTool F1, Makeblock Co., Ltd., Guangdong, China) with specific parameters (200 mm/s speed, 40% power, unidirectional) for LIG patterning[23].
+We utilized a laser system (xTool F1, Makeblock Co., Ltd., Guangdong, China), which is equipped with a 455 nm blue diode laser (rated power: 10 W).
+For LIG fabrication, the laser was operated at 40% power with a scanning speed of 200 mm/s in unidirectional mode [28].
 The LIG sheet was transferred onto a silicone (Ecoflex 00-45 Near Clear, Smooth-On, Inc., Pennsylvania, USA).
@@
 2.3 LIG-transferred soft actuator fabrication process
-Figure 2 illustrates the fabrication process of a soft actuator with the LIG layer positioned adjacent to the liquid interface.
-The fabrication process consists primarily of five steps: (i)–(v).
-Initially, a LIG pattern (2.6 mm $\times$ 40 mm) is created using the aforementioned laser parameters.
-The pattern is then positioned and secured within the mold (detailed in Supplementary Data S1) as shown in (i-a).
-Subsequently, a metallic rod (2.5 mm in diameter) is inserted into the mold, and silicone is introduced and cured, as depicted in (ii).
-In step (iii), the tubular silicone structure with the integrated LIG layer is carefully removed from the mold.
-Because the LIG layer is initially present on the external surface, the tube is inverted in step (iv) to position the LIG layer on the interior.
-In step (v), uncured silicone is introduced into one tube end to create a seal, resulting in an actuator with one open terminus.
-The actuator is completed by introducing 40 µL of a low-boiling-point liquid (Opteon SF33, The Chemours Company, Delaware, USA) into the internal cavity and sealing the open end with a silicone-coated clip.
+Figure 2 illustrates the fabrication of a soft actuator with the LIG layer next to the liquid interface, consisting of five steps (i–v).
+First, a LIG pattern (2.6 mm $\times$ 40 mm) is created with the specified laser parameters.
+Next step, as shown in Fig.
+2(i), the mold and the printed LIG pattern are superimposed and fixed in place.
+Next, a metallic rod (2.5 mm diameter) is placed in the mold, and silicone is poured and cured (Fig.
+2(ii)).
+In Fig.
+2(iii), the tubular silicone structure with integrated LIG is removed.
+Since the LIG is on the exterior, the tube is inverted to position LIG inside (Fig.
+2(iv)).
+In Fig.
+2(v), uncured silicone seals one tube end, yielding an actuator with an open terminus.
+Finally, 40 µL of low-boiling-point liquid (Opteon SF33, Chemours, USA) is added to the cavity, and the end is sealed with a silicone-coated clip.
 Additionally, the silicone sheets (10 mm square with transferred LIG patterns) were fabricated using the same process with molds adjusted to achieve a thickness of 1.0 mm.
-These molds were fabricated using a Formlabs Form 3 printer (Formlabs Inc., Somerville, MA, USA).
+These molds were fabricated using a Formlabs Form 3 printer (Formlabs Inc., MA, USA).
 Clear Resin v4 (Formlabs Inc.) was used as the mold production resin material.
-We examined the temperature response characteristics to investigate whether placing the silicone layer on the exterior and the LIG layer on the interior would affect LIG’s photothermal conversion efficiency.
-A reflector lamp (RF100 V54 WD, 100 V, 54 W; Panasonic Corporation, Osaka, Japan) was used as the light source, and temperature changes on the silicone sheet were measured using a thermal camera (OPTPI64 ILTO15 T090, Optris GmbH, Berlin, Germany).
-Our findings revealed that positioning LIG on the interior actually accelerated the temperature rise at the liquid-contact interface (Supplementary Fig.
+We evaluated temperature response to determine whether placing silicone externally and LIG internally affects photothermal efficiency.
+A reflector lamp (RF100 V54 WD, 100 V, 54 W; Panasonic, Japan) served as the light source.
+Temperature changes on the silicone sheet were measured with a thermal camera (OPTPI64 ILTO15 T090, Optris GmbH,Berlin, Germany), and illuminance was measured using an illuminance meter (LX20, Sanwa Electric Instrument Co., Ltd., Tokyo, Japan).
+Our findings revealed that positioning LIG on the interior accelerated the temperature rise at the liquid-contact interface (Supplementary Fig.
 1).
@@
 2.4 Mechanical and chemical durability of LIG-transferred silicone
-Figure 3 illustrates the silicone sheet (10 mm $\times$ 10 mm) with a transferred LIG layer, as well as the condition of the LIG after being subjected to 20 cycles of stretching and, relaxation, and twisting, respectively.
-In both cases, no delamination of the LIG layer was observed.
-The resistance of the actuator to the liquid was evaluated by immersing the LIG-transferred sheet into a conical tube filled with Opteon SF33.
-The actuator was agitated for 60 seconds.
-After the process, the sheet was examined, and no detachment or degradation of the LIG layer was observed, confirming its stability under these conditions (Fig.
-4).
+To assess the mechanical integrity and interfacial bonding of LIG transferred onto silicone elastomers, we exposed samples to operational stresses.
+The experimental protocol included uniaxial stretching, torsional deformation, and chemical exposure tests using Opteon SF33.
+Test specimens consisted of silicone elastomers with 10 mm $\times$ 10 mm LIG patterns.
+Samples underwent biaxial elongation to 200% strain through 100 loading cycles ( 3A).
+Twisting testing involved $\pm$ 180°rotations repeated over 100 cycles ( 3B).
+Chemical compatibility was examined through 5-minute immersion in Opteon SF33 ( 3C).
+Adhesion was quantified using the tape peel-off test[29, 30].
+After conditioning, 3M Scotch Tape was applied, equilibrated, and then removed from the substrate surface.
+Pre- and post-treatment samples were analyzed by scanning electron microscopy (SEM) to identify delamination.
+Removed tape sections were examined LIG debris.
+Figure 3 presents SEM images of LIG surfaces before and after treatment protocols.
+Imaging was performed using a field-emission SEM (JSM-7800F-PRIME, JEOL Ltd., Tokyo) at 5 kV accelerating voltage.
+Specimens received osmium coating (Tennant 20, Meiwafosis Co., Ltd., Tokyo) prior to examination.
+SEM analysis revealed negligible morphological alterations following stretching, torsional, or immersion treatments.
+The tape peel-off test produced no observable delamination or LIG structural modification, with no detectable transfer residue on tape surfaces.
+These findings demonstrate robust LIG–elastomer interfacial bonding sufficient for anticipated actuator operating conditions.
 3 Heat transfer analysis
@@
 $Q_{hs}=\frac{\sigma(\theta_{h}^{4}-\theta_{s}^{4})\,A}{\dfrac{1}{\epsilon_{h}}+\dfrac{1}{\epsilon_{s}}-1}.$ (1)
-Here, $\sigma=5.67037442\times 10^{-8}~{}\mathrm{W\cdot m^{-2}\cdot K^{-4}}$ is the Stefan-Boltzmann constant; $\theta_{s}$ is the temperature of the silicone wall; $A$ is the heat transfer area; and $\epsilon_{h}$ and $\epsilon_{s}$ are the emissivities of the heat source and the silicone wall, respectively ([24]).
+Here, $\sigma=5.67037442\times 10^{-8}~{}\mathrm{W\cdot m^{-2}\cdot K^{-4}}$ is the Stefan-Boltzmann constant; $\theta_{s}$ is the temperature of the silicone wall; $A$ is the heat transfer area; and $\epsilon_{h}$ and $\epsilon_{s}$ are the emissivities of the heat source and the silicone wall, respectively [31].
 The thermal conductivity of the silicone, $\lambda_{s}$, and its thickness, $x$, are given in Table 1 [32].
@@
 The temperature of the LIG layer, $\theta_{L}(t)$, can be computed by a difference method based on Eqs. (4)–(7).
-The parameters used in the simulation are summarized in Table 1[26].[27]
+The parameters used in the simulation are summarized in Table 1[33, 34].
 Table 1: Parameters used in the simulation
@@
               $h_{se}$               $6~{}\rm{W/m^{2}\cdot K}$
-              $h_{Le}$               $25~{}\rm{W/m^{2}\cdot K}$
+              $h_{Le}$               $18~{}\rm{W/m^{2}\cdot K}$
               $C_{s}$               $1300~{}\rm{J/kg\cdot K}$
@@
               $\alpha_{L}$               $0.83$
-              $Q_{h}$               $0.1~{}\rm{W}$
+              $Q_{h}$               $75~{}\rm{mW}$
 In this simulation, for model simplification, several factors, such as Opteon SF33 vaporization and silicone elastomer elongation, were not taken into account.
@@
 4C.
-When overlaying the temporal temperature variations with the simulation graphs, despite some discrepancies, the T63 values showed close agreement: for the case without LIG, the simulation yielded 118.1 seconds while the experimental measurement was 110.7 seconds; for the case with LIG, the simulation predicted 53.9 seconds compared to the experimental value of 58.3 seconds.
+When overlaying the temporal temperature variations with the simulation graphs, despite some discrepancies, the 63% response time values showed close agreement: for the case without LIG, the simulation yielded 100.0 seconds while the experimental measurement was 103.9 seconds; for the case with LIG, the simulation predicted 54.9 seconds compared to the experimental value of 62.6 seconds.
 These results demonstrate that LIG effectively functions as a photothermal conversion and facilitates heat transfer to the internal liquid.
+The LIG–silicone actuator cooled more rapidly because it reached a higher peak temperature (56.4C vs.
+36.7 C), creating a larger $\delta$ T relative to the ambient ( 25C).
+According to Newton’s law of cooling, this greater temperature difference increases the rate of heat loss.
+Furthermore, the porous LIG structure increases both surface area and thermal conductivity, which also accelerates dissipation.
 4 Evaluation of a soft actuator with transferred LIG under general light source
@@
 5.
-During the experiments, the fabricated actuator was positioned 50 mm away from the light source.
-The light bulb specifications include luminous intensity of 300 cd, total luminous flux of 630 lm, and beam angle of $60^{\circ}$.
-An iPhone 15 Pro Max (Apple Inc., Cupertino, CA, USA) was used to capture the actuator deformation, and ImageJ (citation) was employed for analyzing the bending angles and deformation measurements.
-Figure 6A and C show the bending behavior during light irradiation, the subsequent recovery upon light off, and the corresponding analysis results.
-When exposed to light, the without-LIG actuator exhibited delayed activation, with a 63% response time of 141.7 seconds and approximately 220 seconds required to reach the maximum bending angle.
-In contrast, the proposed actuator with LIG demonstrated a mean 63% response time of 65.0 seconds.
-It reached maximum bending angle in approximately 90 seconds, showing a 54.1% improvement in responsiveness compared to the without-LIG actuator.
-Regarding the recovery phase after light off (Fig 6B and D, the without-LIG actuator showed a 63% response time of 42.7 seconds and required 120 seconds to return to its initial state.
-The proposed actuator with LIG demonstrated enhanced recovery characteristics with a 63% response time of 22.1 seconds and approximately 60 seconds to return to its initial state, representing a 48.2% improvement in recovery performance compared to the without-LIG actuator.
-4.2 Deformation of the actuator before and after light-driven operation
+During experiments, the fabricated actuator was positioned 50 mm from the light source.
+The light bulb had a luminous intensity of 300 cd, a total luminous flux of 630 lm, and a beam angle of 60°.
+An iPhone 15 Pro Max (Apple Inc., Cupertino, CA, USA) captured actuator deformation, and ImageJ analyzed bending angles and deformation measurements[36].
+Figure 6A, C and Supplementary Movie S1 show bending behavior during light irradiation, subsequent recovery upon light off, and corresponding analysis results.
+When exposed to light, the without-LIG actuator exhibited a 63% response time of 141.7 seconds and required approximately 220 seconds to reach maximum bending.
+In contrast, the proposed actuator with LIG demonstrated a mean 63% response time of 65.0 seconds and reached maximum bending in about 90 seconds, a 54.1% improvement.
+During recovery after light off (Fig.
+6, D and Supplementary Movie S1), the without-LIG actuator showed a 63% response time of 42.7 seconds and needed 120 seconds to return to its initial state.
+The proposed actuator with LIG demonstrated a 63% response time of 22.1 seconds and took about 60 seconds to return, a 48.2% improvement in recovery performance.
+To assess repeatability under cyclic light conditions, we performed multiple actuation tests with a single fluid charge (Fig.
+7).
+We encapsulated 40 µL of Opteon SF33 and exposed the actuator to repeated light irradiation.
+The bending angles during the first three cycles were 51.7°, 51.6°, and 50.7°, demonstrating good consistency.
+In the fourth cycle, however, the bending angle declined to 90.4% of the initial peak value and, by the sixth cycle, decreased to 50.5%.
+We attribute this degradation to gas leakage through the silicone elastomer’s gas permeability.
+When we immersed the actuator in water at 65C,, gas bubbles emerged from expanded and thinned areas of the silicone wall (Supplementary Figure 2).
+To address this limitation, we designed the actuator with a refillable configuration and incorporated a detachable clip-based sealing mechanism for repeated operation.
+4.2 Analysis of Actuator Performance Under Different Liquid Volume and Light Distance Conditions
 Figure 8 demonstrates the variations in actuator performance based on changes in the encapsulated liquid volume and the distance between the light source and actuator.
 With the light source fixed at 50 mm, we examined the effect of varying volumes (20 µL, 30 µL, and 40 µL) of encapsulated Opteon SF33.
-The results showed volume-dependent increases in three parameters: the rate of actuator deformation (Fig.
-7A), the plateau bending angle (Fig.
-7B), and the time required to reach the plateau state (Fig.
-7C).
-The linear regression analysis yielded an R² value of approximately 0.98, indicating that the relationship can be effectively approximated as linear within the 20-40 µL range.
-The influence of external environmental conditions was investigated by varying the distance between the light source and the actuator (50 mm, 75 mm, and 10 mm), with results presented in Fig.
-7D-F (Opteon SF33’s amount is 30 µL).
-The deformation ratio and plateau bending angle decreased as the distance increased.
-Notably, the time required to reach the plateau state was the longest in the 100 mm group, showing more than a 2.5-fold increase compared to the 50 mm group.
+The results showed volume-dependent increases in three parameters: the angular change ratio (Fig.
+8A), the plateau bending angle (Fig.
+8B), and the time required to reach the plateau state (Fig.
+8C).
+The linear regression analysis yielded an $R^{2}$ value of approximately 0.98, indicating that the relationship can be effectively approximated as linear within the 20–40 µL range.
+We examined how environmental conditions affect the actuator by testing it at different distances from the light source 50 mm, 75 mm, and 100 mm.
+The results are shown in Fig.
+8D–F using 30µL of Opteon SF33.
+As we moved the actuator further from the light source, both the angular change ratio and the plateau bending angle became smaller.
+The actuator positioned at 100 mm from the light source required a significantly longer time to reach its final bent configuration, approximately 2.5 times greater than that observed at 50 mm distance.
+5 Mechanical characteristics of the actuator
+We evaluated the load-bearing performance of the actuator by attaching polyurethane sponge weights to its distal tip and measuring the resulting reduction in bending angle(Fig.
+9A).
+A nylon thread was connected to the tip of a pre-bent actuator, and sponge-based weights of 0.125, 0.25, 0.5, 0.75, 1.0, and 1.5 g were sequentially suspended via a fine pin.
+The results (Fig.
+9B) indicate that the actuator maintained approximately 90% of its original bending angle under a 0.125 g load ($\approx$ 1.23 mN).
+As the load increased, the actuator’s ability to retain its bent configuration declined, reaching around 50% of the original angle under a 1.5 g load ($\approx$ 14.7 mN).
+These results demonstrate a near-linear relationship between the applied load and bending angle reduction, providing a quantitative evaluation of the actuator’s mechanical output capacity.
 6 Multi-Finger Robot Using Non-Contact & Light-Driven LIG-Transferred Soft Actuators
-To verify the synchronous actuation capability of multiple LIG-transferred soft actuators, we developed a three-fingered robot inspired by the Venus flytrap and evaluated its operation in ambient conditions.
-The demonstration is shown in Movie S2 and Fig.
-8.
-As illustrated in Fig.
-8, upon irradiation from a light source (a lamp in this experiment), the actuators began to bend synchronously, successfully enveloping a ladybug-inspired target object through their coordinated curving motion.
-When the light source was switched off, each actuator was released from its bent state and returned to its original configuration.
+Investigating the practical applications of this actuator, we conducted experimental validation of the multi-finger actuator’s performance in controlled laboratory conditions and assessed its functionality under natural solar illumination.
+6.1 Three-finger actuator performance in laboratory conditions
+To evaluate the feasibility of the proposed actuator system, we investigated synchronized operation of multiple actuators under controlled light exposure conditions and examined the effects of photoactivation and deactivation on actuator performance.
+We assessed the coordinated bending motion of three actuators in response to a light stimulus, aiming to achieve simultaneous coverage of a designated target object.
+The experimental results are documented in Supplementary Movie S2 and Supplementary Figure S3.
+The experimental setup comprised a controlled laboratory environment.
+Illumination was provided by reflector lamps throughout experimental trials.
+The experimental findings demonstrated that upon exposure to incident light at an intensity of approximately 15 klx, all three actuators exhibited synchronized bending behavior with high temporal coordination.
+Upon cessation of light exposure, the actuators consistently returned to their initial linear configuration, confirming the reversible nature of the photoresponsive actuation mechanism.
+6.2 Solar-responsive shading device demonstration
+In natural environments, intense solar radiation can cause harm to flora and fauna.
+To address this challenge, we developed a prototype shading device (Fig.
+10) that responds directly to solar irradiance by providing shade coverage over designated target organisms.
+The device architecture incorporates three actuators integrated with nonwoven fabric.
+We deployed this system outdoors to evaluate solar-driven performance.
+Testing was conducted at 14:00 JST under clear atmospheric conditions.
+On the testing date, the ambient illuminance measured 102.5 klx.
+Using clover as the target organism, we positioned the device under direct solar exposure.
+As demonstrated in Fig.
+10 and Supplementary Movie S3, the actuators responded to sunlight by bending and successfully providing shade coverage over the clover.
+Following deployment, illuminance within the shaded region decreased to 37.6 klx, achieving a 63.3% reduction.
+The actuation process required approximately 140 seconds to complete.
 7 Limitations & future work
@@
 7.2 Gas leakage
-As shown in Fig.
-7F, increasing the distance from the light source decreased the bending angle at a plateau.
-This suggests that Opteon SF33 is leaking out due to silicone’s high gas permeability during the slower actuation caused by reduced light intensity.
-Similar gas leakage phenomena have been discussed in previous studies[15].
-The current actuator design implements a reusable mechanism where one side is sealed with a clip, allowing multiple liquid refilling.
-However, frequent leakage can limit potential applications.
+As shown in Supplementary Fig.
+2, gas leakage phenomena were confirmed to occur in the proposed actuator, similar to previously reported issues[15].
+The current actuator design features a reusable mechanism with one side sealed by a clip, allowing for multiple liquid refilling cycles.
+Future work requires optimal design of the refilling port to enhance sealing performance and improve airtightness.
+However, gas leakage can limit potential applications.
 Potential approaches include incorporating materials with lower gas permeability and developing alternative sealing mechanisms to address this limitation.
 7.3 Utilization of ambient environmental light
-The light source utilized in this study was a reflector lamp (54W) with a beam angle of $60^{\circ}$, center luminous flux of 630 lm, and beam flux of 160 lm.
-Assuming uniform light distribution, calculations based on illuminance indicate that the actuator surface, positioned at a distance of 50 mm, receives approximately 60,000 lux (assuming an actuator LIG surface area of 40 $\rm{mm}^{2}$).
-This illuminance corresponds to approximately 60% of direct summer sunlight[29].
-While outdoor operation would depend on seasonal and geographical factors, implementing lens-based light concentration mechanisms could make outdoor applications feasible.
-Given that typical office environments maintain illuminance levels of approximately 300–1000 lux[30], indoor operation under ambient lighting conditions would be significantly limited.
-Future investigations will explore localized light energy irradiation using laser sources aimed at achieving more precise actuator control at a microscale level.
+The optical characterization and performance evaluation were conducted under controlled laboratory conditions and outdoor environments.
+In the laboratory setup, a reflector lamp (54 W, beam angle: 60°) was employed as the primary light source, exhibiting a center luminous flux of 630 lm and a beam luminous flux of 160 lm.
+Illuminance measurements yielded approximately 15 klx at a working distance of 100 mm and 30 klx at 50 mm, respectively.
+Comparative outdoor experiments under direct solar irradiation demonstrated higher illuminance levels of 100–110 klx, indicating laboratory conditions represented about 30% of natural outdoor intensity.
+The actuator demonstrated slower bending speed outdoors despite increased light intensity, requiring approximately 140 seconds to reach maximum bending angle (Fig.
+10).
+This slower performance stems from convective heat transfer caused by wind, which cools the actuator and prevents the necessary temperature rise.
+Outdoor deployment faces challenges from seasonal changes, location factors, and weather conditions, limiting fast response times.
+Optical concentration systems, such as focusing lenses, could boost light energy density on the LIG surface.
+Smart use of wind-induced cooling might enable quicker bending–unbending cycles, broadening outdoor applicability.
+Additionally, localized light energy irradiation using laser sources to improve actuator control precision at the microscale represents a promising avenue for future research.
 8 Conclusions
-This paper focused on the ocular structure of nocturnal animals that efficiently collect and utilize light in low-irradiation environments.
 We addressed the limitation of conventional silicone-based soft actuators, which exhibited low driving efficiency due to ineffective light utilization.
-We incorporated a light-heat conversion element layer called LIG on the light-incident surface.
+We incorporated a photothermal conversion layer using LIG on the light-incident surface.
 This modification enabled the actuator to effectively utilize light that would otherwise be poorly absorbed by silicone alone, heating the internal low-boiling-point liquid.
 As a result, the proposed actuator with LIG demonstrated a 54% improvement in response time compared to conventional without-LIG actuators.
-While various approaches have successfully enhanced photothermal conversion and thermal conductivity by incorporating graphene or metal powders into silicone elastomers, our method explores an alternative approach by installing a LIG layer on the silicone surface.
-This approach preserves the inherent properties of silicone while utilizing LIG, which can be manufactured under standard environmental conditions, thereby facilitating both the production of photothermal actuators and their diverse applications.
-Future research will focus on the versatility of LIG transfer patterns, aiming to develop efficient light-utilizing LIG patterns under natural environmental conditions and design more complex actuator geometries with corresponding LIG patterns.
-This research direction aligns with the principles of soft robotics, pursuing the development of soft actuators with enhanced deformation capabilities and environmental adaptability.
+Our LIG-based approach preserves silicone’s inherent characteristics while enabling fabrication under ambient conditions and operation with environmental light sources including sunlight.
+This method significantly simplifies the manufacturing process of photothermal actuators and opens up promising possibilities for their broader application in diverse fields.
 Acknowledgement
-This work was supported by JSPS KAKENHI Grant Number JP24K21213.
+This work was supported by the Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Number JP24K21213.
+This work was supported by "Advanced Research Infrastructure for Materials and Nanotechnology in Japan(ARIM)" of the Ministry of Education, Culture, Sports, Science and Technology (MEXT), Grant Number JPMXP1225UT0165.
 M.S.
-sincerely thank my cats, whose eyes served as a visual reference for the demonstration of the tapetum lucidum structure and whose presence inspired and supported me throughout the research process.
+sincerely thanks T.
+Kondo, who kindly provided support in the use of the scanning electron microscope, and my cats, whose eyes served as a visual reference for demonstrating the tapetum lucidum structure and whose presence inspired and supported me throughout the research process.
 Author declaration
@@
   1. 1.
-    Supplementary Figure S1
+    Supplementary Figure S1:Effect of light irradiation direction Temperature changes on the opposite surface (liquid-contact surface) when the light was irradiated from one direction onto either the silicone side (A) or LIG side (B) of silicone elastomer with transferred LIG or silicone only (C).
+The fastest temperature rise was observed when light was irradiated from the silicone side (case A).
   2. 2.
-    Supplementary Movie The actuator’s bending under light irradiation and its bending release upon turning off the light.
+    Supplementary Figure S2: Visualization of gas leakage from the proposed actuator.
+The proposed actuator was immersed in 65C hot water to visualize gas leakage from within.
+Small bubbles can be seen leaking from the expanded actuator.
   3. 3.
-    Supplementary Movie S2:Demonstration of a bio-inspired (Venus flytrap) robot utilizing the proposed actuator in a biomimetic movement driven by light.
+    Supplementary Figure S3: Synchronized actuation of three-finger actuator system under light control.
+A: Schematic illustration showing the operating principle: actuators remain straight when light is OFF and bend to form a protective enclosure when light is ON through a "hold and protect" mechanism.
+B: Time-lapse demonstration of synchronized actuation in laboratory conditions showing three actuators simultaneously bending to cover a target object upon light exposure (15 klx).
+At t = 0.00 seconds (light-off), the actuators maintain a straight configuration.
+At t = 1.00 seconds (light-on), the actuators exhibit a coordinated bending motion, forming a protective enclosure.
+At t = 2.20 seconds (light-off), the actuators return to their initial straight state, demonstrating reversible and controllable actuation through light activation and deactivation.
+  4. 4.
+    Supplementary Movie S1: The actuator’s bending under light irradiation and its bending release upon turning off the light (related to Fig. 6).
+  5. 5.
+    Supplementary Movie S2: Demonstration of a multi-finger robot in a laboratory setup mimicking a heterogeneous natural environment.
+The ambient temperature was 20C, and the illuminance was 30.2 klx.
+Under light irradiation, the three actuators moved synchronously, and upon light-off, their bending was released (see Supplementary Figure S3).
+  6. 6.
+    Supplementary Movie S3: Demonstration of shading robot driven by sunlight in nature (related to Fig. 10).